### v4

# Transductions of Graph Classes Admitting Product Structure

Abstract
In a quest to thoroughly understand the first-order transduction hierarchy of hereditary graph classes, some questions in particular stand out; such as, what properties hold for graph classes that are first-order transductions of planar graphs (and of similar classes)? When addressing this (so-far wide open) question, we turn to the concept of a product structure – being a subgraph of the strong product of a path and a graph of bounded tree-width, introduced by Dujmović et al. [JACM 2020]. Namely, we prove that any graph class which is a first-order transduction of a class admitting such product structure, up to perturbations also meets a structural description generalizing the concept of a product structure in a dense hereditary way — the latter concept being introduced just recently by authors under the name of ${\cal H}$-clique-width [MFCS 2024].
Using this characterization, we show that the class of the 3D grids, as well as a class of certain modifications of 2D grids, are not first-order transducible from classes admitting a product structure, and in particular not from the class of planar graphs.
keywords:
product structure, hereditary class, clique-width, twin-width
1 Introduction
A recent research trend on the boundary between structural graph theory and graph logic is to study first-order (FO) transductions between (hereditary) graph classes, e.g. [DBLP:journals/corr/abs-2208-14412, DBLP:journals/corr/abs-2501.04166]. This research is motivated by structural questions about dense hereditary graph classes, and by the significant goal to resolve computational complexity of the FO model checking problem on dense graph classes.
We refer to Section 2 for the definitions. Briefly, a simple first-order interpretation is given by a binary FO formula $\xi$ over graphs, and for a graph $G$ the result of the interpretation is the graph $\xi(G)$ on the same vertex set and the edge relation (assumed symmetric) determined by $uv\in E(\xi(G))$ $\iff$ $G\models\xi(u,v)$. In a first-order transduction $\tau$, one can additionally duplicate the vertices of $G$ and assign arbitrary vertex colors before applying $\xi$, and then take an induced subgraph of the result (in particular, $\tau(G)$ is actually a set of graphs – unlike $\xi(G)$). These notions are naturally extended to graph classes; here $\tau({\cal C})$ denotes the class of all graphs obtained by applying $\tau$ to (colored) graphs from a class ${\cal C}$.
All transductions in this paper are first-order. We say that a graph class ${\cal D}$ is transducible from a class ${\cal C}$ if there is a transduction $\tau$ such that ${\cal D}\subseteq\tau({\cal C})$. We interpret this relation as that ${\cal D}$ is not “logically richer” than ${\cal C}$ (although ${\cal D}$ may be combinatorially a lot more complicated class than ${\cal C}$), and the aim is that if we understand and can handle the class ${\cal C}$, we would be able to do the same with ${\cal D}$.
There are two core and mutually related questions in the area; for a class ${\cal C}$, describe structural properties satisfied by (all) transductions of ${\cal C}$, and for classes ${\cal C},{\cal D}$, decide whether ${\cal D}$ is transducible from ${\cal C}$. For an illustration, let ${\cal T}$ be the class of trees, then all transductions of ${\cal T}$ are of bounded clique-width, but not every graph class of bounded clique-width is transducible from ${\cal T}$ (this would be true considering richer MSO transductions). In fact, Nešetřil et al. [DBLP:conf/soda/NesetrilMPRS21] showed that a graph class ${\cal C}$ is a first-order transduction of a class of graphs of bounded tree-width, if and only if ${\cal C}$ is of bounded clique-width and excludes some half-graph as a bi-induced subgraph (we say that it has stable edge relation).
Results of this kind can be useful also in the FO model checking problem. Consider for example the following result of Gajarský et al. [DBLP:journals/tocl/GajarskyHOLR20]; a graph class ${\cal D}$ is a transduction of a graph class of bounded maximum degree, if and only if ${\cal D}$ is near-uniform. Since, luckily, the claimed transduction can be efficiently computed, the characterization [DBLP:journals/tocl/GajarskyHOLR20] readily gives an efficient FO model checking algorithm for each near-uniform graph class. In general, when a graph class ${\cal C}$ has an efficient FO model checking algorithm, and a class ${\cal D}$ is transducible from ${\cal C}$ (say, via $\tau$), there are still two obstacles for FO model checking on ${\cal D}$ : First, for $G\in{\cal D}$ one has to be able to compute $G_{1}\in{\cal C}$ such that $G\in\tau(G_{1})$, and second, one also has to guess the coloring of $G_{1}$ used in $\tau$ to obtain $G$.
On the other hand, it is also interesting and useful to understand why a graph class ${\cal D}$ is not transducible from a graph class ${\cal C}$. While on the positive side (‘is transducible’), one can simply give a transduction and guess a coloring, on the negative side the task is much more difficult—one has to be able to find, for every transduction $\tau$, a graph in ${\cal D}\setminus\tau({\cal C})$ and that is hard to argue. In other words, in ${\cal D}$, one needs to understand the ‘obstacles’ which are not transducible from ${\cal C}$.
For example, long paths are such obstructions for transducibility from a class of bounded-height trees. This is in fact a tight characterization; Ossona de Mendez, Pilipczuk and Siebertz [DBLP:journals/ejc/MendezPS25] proved that a graph class ${\cal C}$ is of bounded shrub-depth (that is, transducible from a class of bounded-height trees), if and only if ${\cal C}$ does not transduce the class of all paths. Other, more complex examples of how understanding the obstructions for transducibility is used to obtain strong results are contained, e.g., in Bonnet et al. [DBLP:journals/jacm/BonnetGMSTT24] or in Dreier, Mählmann and Torunczyk [DBLP:conf/stoc/DreierMT24].
However, except sporadic examples of success, we currently lack good tools to prove non-transducibility. A typical approach (to proving that a class ${\cal D}$ is not transducible from ${\cal C}$) suggests a structural property $\Pi$ satisfied by ${\cal C}$ and preserved under transductions, and then finds graphs in ${\cal D}$ not satisfying $\Pi$. The rather few examples of properties preserved under transductions include, e.g.; monadic stability [SHELAHstabil], bounded clique-width [DBLP:journals/mst/CourcelleMR00], each value of shrub-depth [DBLP:journals/lmcs/GanianHNOM19], bounded twin-width [DBLP:journals/jacm/BonnetKTW22] and bounded flip-width [DBLP:conf/focs/Torunczyk23].
1.1 Our results
Our research is inspired by the task to characterize classes which are transducible from the class of planar graphs. While planar graphs are usually perceived as combinatorially and algorithmically “nice”, their transductions (including, e.g., the well-studied classes of $k$-planar and fan-planar graphs, or map graphs) can be combinatorially very complicated. (For example, testing 1-planarity is NP-hard already for planar graphs with one added edge [DBLP:journals/siamcomp/CabelloM13].) It is an open question whether the class of toroidal graphs (embeddable on the torus) is transducible from planar graphs.
Not much was known prior to this paper about obstructions to transducibility from planar graphs except that the class of planar graphs is monadically stable and, more recently, that such transductions are of bounded twin-width [DBLP:journals/jacm/BonnetKTW22]. One can also prove, using known logical tools and the fact that planar graphs are of bounded local tree-width, that classes transducible from planar graphs are of bounded local clique-width up to perturbations.
On the other hand, several major breakthrough results about planar graphs in the combinatorics domain, such as getting an upper bound on the queue number of planar graphs, have been recently obtained using a new tool of a product structure [DBLP:journals/jacm/DujmovicJMMUW20]. In the original basic setting, a class ${\cal C}$ admits a product structure if, for some $k$, every graph in ${\cal C}$ is a subgraph of the strong product $\boxtimes$ (Figure 1) of a path and a graph of tree-width $\leq k$. See Definition 2.1. Planar and surface-embeddable graphs, for instance, do admit product structure (Theorem 2.2).
Our main result claims that transductions of classes admitting a product structure essentially (up to bounded perturbations) follow a similar kind of product structure (here briefly reformulated in view of Definition 2.3 and Theorem 2.4):
  • (Theorem 3.19) If ${\cal C}$ is a class admitting a product structure and $\tau$ is a transduction defining the class of simple graphs $\tau({\cal C})$, then every graph in $\tau({\cal C})$ is a bounded perturbation of an induced subgraph of the strong product of a path and a graph from a class of bounded clique-width.
We actually prove, in Theorem 3.16, a generalized version in which paths (of the assumed product structure) are replaced by graphs from any fixed class of bounded maximum degree.
We also give the converse direction of the main result, in a modification consisting in strengthening the assumption of bounded clique-width to bounded stable clique-width:
  • (Theorem 4.2) Assume that every graph in a class ${\cal C}$ is a bounded perturbation of an induced subgraph of the strong product of a path and a graph from a class of bounded stable clique-width. Then, ${\cal C}$ is transducible from a class admitting a product structure.
Again, this is also generalized to any fixed class of bounded maximum degree in place of the class of paths.
Finally, we demonstrate the strength of Theorem 3.19 by giving straightforward proofs of the following consequences:
  1. I.
    (Corollary 5.7) The class of all 3D grids, and
  2. II.
    (Corollary 5.13) the class of certain non-local modifications of 2D grids – by adding a bunch of apex vertices adjacent to mutually distant nodes of the grid,
are not transducible from any class admitting product structure, and in particular not from the classes of planar graphs and graphs embeddable on a fixed surface.
At the same time and independently of us, Gajarský, Pilipczuk and Pokrývka proved point I. in [DBLP:journals/corr/abs-2501-07558].
1.2 Proof outline
The core of our results is in the following technical lemma:
  • (Lemma 3.1) Let $Q$ be a graph of maximum degree $d$, $Q^{\circ}_{r}$ denote the reflexive $r$-th power of $Q$, and $M$ be a graph of tree-width $k$. Every strongly $r$-local interpretation $\xi$ of any 2-colored spanning subgraph of the product $Q\boxtimes M$ (cf. Definition 2.1) has $\{Q^{\circ}_{r}\}$-clique-width (Definition 2.3) bounded in terms of $r,d,k$ and the rank of $\xi$.
We note that, the requirement of ${\cal Q}$ having bounded degree is necessary and best possible, since Hliněný and Jedelský [DBLP:conf/mfcs/HlinenyJ24] (in the extended arXiv version) proved that already the class of strong products of two stars is monadically independent, that is, by suitably coloring it, one can transduce all graphs from it.
We informally describe the proof idea of Lemma 3.1:
Let $G\subseteq Q\boxtimes M$, where $Q$ and $M$ are as above, such that $V(G)=V(Q\boxtimes M)$. Let $\xi$ be a strongly local formula. The idea of the proof of Lemma 3.1 is to recursively, in a bottom-up fashion, build a $(Q,f(\ldots))$-expression (Definition 2.3) for the subgraphs $G^{\prime}$ of $\xi(G)$ induced by $V(Q\boxtimes M^{\prime})$, where $M^{\prime}\subseteq M$ runs over subgraphs defined by subtrees in the tree decomposition of $M$. The color of each vertex $v$ of $G^{\prime}$ in the expression encodes (via Definition 3.5) its type to all vertices outside $G^{\prime}$ that are reachable from $v$ by a path of bounded length.
Using the information stored in the vertex colors and $r$-locality of $\xi$, we decide the existence of an edge $uv$ in $\xi(G)$ depending only on the colors of $u$ and $v$ and intersection of $r$-balls around $u$ and $v$. We prove correctness of this decision in Claims 4 and 5. Thus, the color of each vertex $v$ of $G^{\prime}$ encodes its $r$-neighborhood outside $G^{\prime}$. Encoding the $r$-neighborhoods naively would result in an unbounded number of colors. There are two core ideas enabling us to reuse colors (as embedded in Definition 3.5 and wrapped-up by 12):
  • Strong locality of $\xi$ guarantees that we only ever create edges between vertices with intersecting $r$-balls, and so we can use the parameter graph $Q^{\circ}_{r}$ to avoid creating edges at large $Q$-distance in the product $Q\boxtimes M$. Consequently, we may “remember” the vertices of $Q$ by only their colors in a proper coloring of the $3r$-power of $Q$.
  • We do not really remember all vertices of $V(G)$ in the intersection of the $r$-neighborhoods of every pair of already-created vertices $u$ and $v$ of $G^{\prime}$. In fact, it suffices to remember the vertices of some $u$$v$ separator of the union of the $r^{\prime}$-neighborhoods of $u$ and of $v$ for some $r^{\prime}$ depending on $r$ and the rank of $\xi$. We use weak colorings to bound the number of vertices $m\in V(M)\setminus V(M^{\prime})$ such that we have to encode some information concerning them in the color of a vertex $v$ of $V(G^{\prime})$.
1.3 Paper structure
Section 2 surveys the main related concepts and definitions. Section 3 proves Lemma 3.1 and concludes with the main results, Theorem 3.16 and Theorem 3.19. Section 4 shows the stable converse direction of the main results, and Section 5 proves the mentioned consequences on non-transducibility of certain grids. Section 6 adds concluding remarks.
2 Preliminaries
2.1 Graphs, decompositions, and weak colorings
We write $[k]$ to denote the set $\{1,2,\ldots,k\}$. In this paper, unless stated otherwise, graphs are simple and finite, loop graphs additionally allow loops, and reflexive graphs are those with a loop incident to every vertex. Let $V(G)$ denote the vertex set of a graph $G$ and $E(G)$ denote its edge set, and let $G\subseteq H$ mean that $G$ is a subgraph of $H$ and $G\subseteq_{i}H$ that $G$ is an induced subgraph of $H$. A graph class ${\cal C}$ is hereditary if ${\cal C}$ is closed under taking induced subgraphs.
Let an $a\times b$-grid ($a\times b\times c$-grid) be the graph which is the Cartesian product [^1] Recall that in the Cartesian product of graphs, two tuples are adjacent iff they agree in all coordinates except one in which they form an edge (of the respective factor). of two paths on $a$ and $b$ vertices (three paths on $a$ and $b$ and $c$ vertices). We also shortly call these graphs 2D and 3D grids.
For an integer $k\geq 1$, let the $k$-th power of a graph $G$ be the graph on the same vertex set and edges joining exactly the pairs of distinct vertices which are at distance at most $k$ in $G$.
A graph $G$ is a half-graph of order $n$ if $G$ is a bipartite graph with the bipartition $\{u_{1},\ldots,u_{n}\}$ and $\{v_{1},\ldots,v_{n}\}$, such that $u_{i}v_{j}\in E(G)$ if and only if $i\leq j$. A bipartite graph $G$ with a fixed bipartition $V(G)=A\cup B$ is a bi-induced subgraph of a graph $H$, if $G\subseteq H$ such that every edge of $H$ with one end in $A$ and the other end in $B$ is present in $G$.
A tree decomposition of a graph $G$ is a pair $(T,{\cal X})$, where $T$ is a tree and ${\cal X}=(X_{t})_{t\in T}$ is a collection of bags $X_{t}\subseteq V(G)$ indexed by the nodes of $T$ satisfying the following properties:
  • $\bigcup_{t\in V(T)}X_{t}=V(G)$
  • For every edge $uv\in E(G)$, there is a bag $X_{t}$ where $t\in V(T)$ containing both $u$ and $v$.
  • For every vertex $v\in V(G)$, the set of the nodes $t$ of $T$ such that $v\in X_{t}$ induces a subtree of $T$. This is called the interpolation property.
Width of a tree decomposition $(T,{\cal X})$ is $\max_{t\in V(T)}|X_{t}|-1$. The tree-width of a graph $G$ is the minimum width over all tree decompositions of $G$. We say that a tree decomposition $(T,{\cal X})$ is smooth if, for every edge $t_{1}t_{2}\in E(T)$ of the decomposition tree $T$, we have that ${|X_{t_{1}}\setminus X_{t_{2}}|\leq 1}$ and ${|X_{t_{2}}\setminus X_{t_{1}}|\leq 1}$. Observe that every tree decomposition can be modified to become smooth without increasing its width.
A graph class ${\cal C}$ is said to be of bounded local tree-width if there exists an integer function $g$ such that, for every $G\in{\cal C}$, $v\in V(G)$ and every $r$, the tree-width of the subgraph induced by the neighborhood of $v$ in $G$ up to distance $r$ is at most $g(r)$.
Given a pair of graphs $G$ and $H$, their strong product $G\boxtimes H$ is a graph on the vertex set $V(G\boxtimes H)=V(G)\times V(H)=\{[g,h]|\>g\in V(G),h\in V(H)\}$ such that $\{[g_{1},h_{1}],[g_{2},h_{2}]\}$ is an edge of $G\boxtimes H$ if and only if one of the following is true;
  • $g_{1}=g_{2}$ and $h_{1}h_{2}\in E(H)$, or
  • $g_{1}g_{2}\in E(G)$ and $h_{1}=h_{2}$, or
  • $g_{1}g_{2}\in E(G)$ and $h_{1}h_{2}\in E(H)$.
See Figure 1 for an illustration. Notice that throughout the paper we use to refer to the vertices of the product $G\boxtimes H$ as to pairs $[g,h]$ where $g\in V(G)$ and $h\in V(H)$.
Figure 1: Illustrating the strong product $\boxtimes$ of the two shaded graphs.
Definition 2.1 (Product structure).
For any class ${\cal Q}$ of bounded-degree graphs, we say that a graph class ${\cal C}$ admits ${\cal Q}$-product structure if there is a constant $k$ (depending on ${\cal C}$), such that for every graph $G\in{\cal C}$ there is a graph $Q\in{\cal Q}$ and a graph $M$ of tree-width at most $k$ such that $G\subseteq Q\boxtimes M$. Specially, we say that a graph class ${\cal C}$ admits product structure if ${\cal C}$ admits ${\cal Q}$-product structure where ${\cal Q}$ is the class of paths. (In the latter case, we sometimes add an adjective classical to such product structure with paths.)
Notably, the following hold:
Theorem 2.2 ([DBLP:journals/jacm/DujmovicJMMUW20, DBLP:journals/combinatorics/UeckerdtWY22, DBLP:journals/dmtcs/DistelHHW22, DBLP:journals/jctb/DujmovicMW23, DBLP:journals/corr/abs-2001-08860]).
Graph classes admitting (the classical) product structure include, e.g.:
  • planar graphs, and graphs embeddable on a fixed surface,
  • graphs drawable in the plane or a fixed surface with bounded number of crossings per edge ($k$-planar graphs),
  • apex-minor-free graphs.
Moreover, a mild extension of product structure (allowing for apex vertices and clique-sums) holds for every proper minor-closed class [DBLP:journals/jacm/DujmovicJMMUW20].
Finally, we recall the notion of weak coloring numbers. Let $G$ be a graph and let $\preceq_{G}$ be a linear order of its vertex set $V(G)$. We say that a vertex $u$ is weakly $r$-reachable (with respect to $\preceq_{G}$) from a vertex $v$ if there is a $u$$v$ path $P$ of length at most $r$ such that, for all vertices $w\in V(P)$, we have $u\preceq_{G}w$. We denote by ${\text{WReach}_{r}[G,\preceq_{G},v]}$ the set of all vertices weakly $r$-reachable from $u$. The weak coloring number of a graph $G$ is the minimum over all linear orders $\preceq_{G}$ of $\max_{v\in V(G)}\big{|}{\text{WReach}_{r}[G,\preceq_{G},v]}\big{|}$.
2.2 ${\cal H}$-clique-width
We use the following new concept of Hliněný and Jedelský [DBLP:conf/mfcs/HlinenyJ24] extending the traditional clique-width structural measure [DBLP:conf/gg/CourcelleER90]. Let $H$ be a loop graph and let $k\geq 2$ be an integer. Consider labeled graphs, where each vertex $v$ is assigned a label $(h,c)$ such that $h\in V(H)$ is the parameter vertex of $v$ and $c\in[k]$ is the color of $v$.
Definition 2.3.
An $(H,k)$-expression is an algebraic expression composed of the following operations whose values are labeled graphs as considered above:
  1. a)
    Given a parameter vertex $h\in V(H)$ and a color $c\in[k]$, create a one-vertex graph with the vertex labeled $(h,c)$.
  2. b)
    Take the disjoint union of two labeled graphs.
  3. c)
    Given a pair of colors $c_{1},c_{2}\in[k]$, recolor $c_{1}$ to $c_{2}$ in a labeled graph; that is, replace each label $(h,c_{1})$ by label $(h,c_{2})$. Notably, this operation only modifies the colors, it does not modify the parameter vertices.
  4. d)
    Given a pair of colors $c_{1},c_{2}\in[k]$, add edges between colors $c_{1}$ and $c_{2}$; that is, make adjacent all pairs of vertices $u$ and $v$ such that the label of $u$ is $(h,c_{1})$, the label of $v$ is $(h^{\prime},c_{2})$, and the parameter vertices of $u$ and $v$ are adjacent in $H$; $hh^{\prime}\in E(H)$.
We refer to Figure 2 for an illustration of Definition 2.3, and to [DBLP:conf/mfcs/HlinenyJ24] for combinatorial properties of this notion. In particular, observe that the mapping from the vertices of an expression to their parameter vertices is a graph homomorphism to $H$.
[…]
$p_{4}$$p_{3}$$p_{2}$$p_{1}$$p_{4}$$p_{2}$$p_{3}$$p_{1}$$p_{4}$$p_{2}$$p_{3}$$p_{1}$(f)$p_{4}$$p_{2}$$p_{3}$$p_{1}$$p_{4}$$p_{2}$$p_{3}$$p_{1}$(g)$p_{4}$$p_{2}$$p_{3}$$p_{1}$$p_{4}$$p_{2}$$p_{3}$$p_{1}$(h)
$p_{4}$$p_{3}$$p_{2}$$p_{1}$$p_{4}$$p_{2}$$p_{3}$$p_{1}$$p_{4}$$p_{2}$$p_{3}$$p_{1}$$p_{4}$$p_{2}$$p_{3}$$p_{1}$(i)$p_{4}$$p_{2}$$p_{3}$$p_{1}$$p_{4}$$p_{2}$$p_{3}$$p_{1}$$p_{4}$$p_{2}$$p_{3}$$p_{1}$(j)$p_{4}$$p_{2}$$p_{3}$$p_{1}$$p_{4}$$p_{2}$$p_{3}$$p_{1}$$p_{4}$$p_{2}$$p_{3}$$p_{1}$$p_{4}$$p_{2}$$p_{3}$$p_{1}$(k)
$p_{4}$$p_{3}$$p_{2}$$p_{1}$$p_{4}$$p_{2}$$p_{3}$$p_{1}$$p_{4}$$p_{2}$$p_{3}$$p_{1}$$p_{4}$$p_{2}$$p_{3}$$p_{1}$$p_{4}$$p_{2}$$p_{3}$$p_{1}$$p_{4}$$p_{2}$$p_{3}$$p_{1}$$p_{4}$$p_{2}$$p_{3}$$p_{1}$$p_{4}$$p_{2}$$p_{3}$$p_{1}$(l)
Figure 2: An informal illustration of a $(P,5)$-expression making the $a\times b\,$-grid, where $P$ is a $b$-vertex reflexive path (here $b=4$). Each vertex in the expression is labeled $(p_{i},c)$ where $p_{i}\in V(P)$ is the parameter vertex inscribed in the vertex, and $c\in[5]$ is the color, here depicted as light/dark red or green, or blue. In the first phase, see (a)–(e), we create the vertical edges of the grid, independently for each column (with alternating shades of red or green). In the second phase, see (f)–(l), we union the columns and add by the definition precisely the required edges in the rows – between dark red and dark green, and separately between light red and light green colors.
Given a (fixed) class of loop graphs ${\cal H}$, we define the ${\cal H}$-clique-width of a (simple) graph $G$ as the smallest integer $\ell$ such that (a coloring of) the graph $G$ is the value of an $(H,\ell)$-expression for some $H\in{\cal H}$.
The concept is closely related to the aforementioned product structure via the following rather straightforward statement:
Theorem 2.4 (Hliněný and Jedelský [DBLP:conf/mfcs/HlinenyJ24]).
Let ${\cal H}$ be a family of reflexive graphs and ${\cal H}^{\prime}$ be the family obtained from ${\cal H}$ by removing all loops of all graphs. For every $\ell\geq 2$; a simple graph $G$ is of ${\cal H}$-clique-width at most $\ell$, if and only if $G$ is an induced subgraph of the strong product $H^{\prime}\boxtimes M$ where $H^{\prime}\in{\cal H}^{\prime}$ and $M$ is a simple graph of clique-width at most $\ell$.
Considering an $(H,\ell)$-expression $\phi$, the deparameterization of $\phi$ is an expression $\phi_{1}$ obtained as follows: Let $K_{1}^{\circ}$ denote a single vertex graph formed by a vertex $v$ with a loop. Then $\phi_{1}$ is the $(K_{1}^{\circ},\ell)$-expression constructed from $\phi$ by replacing every parameter vertex with $v$. (Note that “deparameterized ${\cal H}$-clique-width” is exactly the usual clique-width.) Furthermore, we say that an $(H,\ell)$-expression $\phi$ is $k$-stable if the value of the deparameterization $\phi_{1}$ of $\phi$ does not contain a half-graph of order $k$ as a bi-induced subgraph.
2.3 Relational structures and logic
A relational structure is an $\ell$-tuple $(V,R_{1},\ldots,R_{\ell})$, where $V$ is the universe and $R_{i}\subseteq V^{a_{i}}$ is a relation of arity $a_{i}$. The signature of such relational structure is the tuple $(R_{1},\ldots,R_{\ell})$ of symbols together with their arities. We view $c$-colored graphs $G=(V,E,C_{1},\ldots,C_{c})$ as relational structures over a signature with the binary edge relation and unary relations for each color.
In this article, unless we say otherwise, we only consider first-order logic, thus all our formulas are first-order. In order to simplify our notation, we assume that free variables of each formula are in the set $\{x_{i}|i\in\mathbb{N}\}$ and bound variables are in the set $\{y_{i}|i\in\mathbb{N}\}$. We write $\phi(x_{1},x_{2},\ldots,x_{k})$ to denote that the set of free variables of $\phi$ is a subset of $\{x_{1},x_{2},\ldots,x_{k}\}$. Then, given a relational structure $G$ and a $k$-tuple $\bar{v}=(v_{1},v_{2},\ldots,v_{k})$ of elements of $V(G)$, we write $G\models\phi(\bar{v})$ to denote $G,(x_{1}\mapsto v_{1}),(x_{2}\mapsto v_{2}),\ldots,(x_{k}\mapsto v_{k})\models\phi$, where $x_{k}\mapsto v_{k}$ denotes that the value of $x_{k}$ is $v_{k}$.
We often simply say that a formula $\phi$ has arity $k$ to denote $\phi(x_{1},x_{2},\ldots,x_{k})$. Given a formula $\phi(x_{1},\ldots,x_{k})$ and indices $i_{1},i_{2},\ldots,i_{k}$, we denote by $\phi[x_{1}\to x_{i_{1}},x_{2}\to x_{i_{2}},\ldots,x_{k}\to x_{i_{k}}]$ the formula obtained from $\phi$ by renaming $x_{j}$ to $x_{i_{j}}$ for all $1\leq j\leq k$. We note that, due to the assumption from the beginning of this paragraph, the renaming only affects free variables, it does not affect any bound variable.
Given a signature $\sigma=(R_{1},\ldots,R_{\ell})$ and formulas $\xi_{1},\ldots,\xi_{\ell^{\prime}}$ together with their arities $a_{1},\ldots,a_{\ell^{\prime}}$ and a formula $\omega$ of arity $1$, using predicates from $\sigma$, the $(\xi_{1},\ldots,\xi_{\ell^{\prime}},\>\omega)$-interpretation of a relational structure $G$ over $\sigma$ is the relational structure $H=(V^{H},R_{1}^{H},\ldots,R_{\ell^{\prime}}^{H})$, where $v\in V^{H}\subseteq V^{G}$ iff $G\models\omega(v)$, and each $R_{i}^{H}$ has the same arity $a_{i}$ as $\xi_{i}$ and $(v_{1},\ldots,v_{a_{i}})\in R_{i}^{H}$ iff $G\models\xi_{i}$. An interpretation is first-order if all involved formulas are first-order.
Given a formula $\xi$ and a relational structure $G$, we denote by $\xi(G)$ the result of the $(\xi,\top)$-interpretation of $G$, where $\top$ is just tautology (of arity 1). A $k$-copy of a relational structure $G=(V^{G},R_{1}^{G},\ldots,R_{\ell}^{G})$ is the relational structure $H=(V^{G}\times[k],R_{1}^{H},\ldots,R_{\ell}^{H},\>SC,SV)$ obtained from $G$ by taking the disjoint union of $k$ copies of $G$ and adding binary relations $SV=\{([u,i],[u,j])|\>u\in V^{G},i,j\in[k]\}$ and $SC=\{([u,i],[v,i])|\>u,v\in V^{G},i\in[k]\}$. A $k$-coloring of a relational structure $G=(V^{G},R_{1}^{G},\ldots,R_{\ell}^{G})$ is the operation resulting in the class of relational structures $H=(V^{G},$ $R_{1}^{G},\ldots,R_{\ell}^{G},$ $C_{1},\ldots,C_{k})$ where $C_{i}$ are interpreted as arbitrary unary relations over $V^{G}$.
The operations of interpretation, copying, and coloring naturally extend to classes of relational structures by applying them to each member of the class. A (first-order) transduction is any composition of copying, coloring, and a (first-order) interpretation. Given a class ${\cal C}$ and a transduction $\tau$, we denote by $\tau({\cal C})$ the class obtained from ${\cal C}$ by applying $\tau$. In this paper, unless stated otherwise, we consider transductions of classes of colored graphs producing classes of graphs.
A transduction is non-copying if the copying operation is not used. We say that a class ${\cal C}$ does not need copying if, for each $k$, there is a non-copying transduction $\tau_{k}$ such that $k{\cal C}\subseteq\tau_{k}({\cal C})$, where $k{\cal C}$ denotes the $k$-copy of ${\cal C}$. Similarly, we say that ${\cal C}$ does not need coloring [^2] See 13 for an example of such class. if, for each $k$, there is an interpretation $\xi_{k}$ such that ${\cal C}^{c}\subseteq\xi_{k}({\cal C})$ for every $k$-coloring ${\cal C}^{c}$ of ${\cal C}$.
[…]
We say that a non-copying transduction is immersive if the formulas defining the involved interpretations are strongly local. We say that a transduction $\tau$ is subsumed by a transduction $\upsilon$ if for every class ${\cal C}$ we get $\tau({\cal C})\subseteq\upsilon({\cal C})$.
We say that a graph $G$ is a $k$-perturbation of a graph $H$ if $V(G)=V(H)$ and there is a partition ${\cal P}=(P_{1},P_{2},\ldots P_{k})$ of $V(G)=V(H)$ into $k$ parts and a symmetric relation $\sim$ on $[k]$ such that the following holds: For every pair of distinct vertices $u\in P_{i}$ and $v\in P_{j}$ where $i,j\in[k]$ (possibly $i=j$), $uv\in E(G)$ if and only if $(uv\in E(H)\iff i\sim j)$. The notion naturally extends to graph classes; a class ${\cal C}$ is a $k$-perturbation of a class ${\cal D}$ if each graph of ${\cal C}$ is a $k$-perturbation of some graph of ${\cal D}$. Perturbations are sometimes also called flips in the literature.
Theorem 2.5 (Nešetřil, Ossona de Mendez, and Siebertz [DBLP:conf/csl/NesetrilMS22]).
Any non-copying first-order transduction is subsumed by the composition of an immersive transduction and a perturbation.
In the coming proof, we employ the notion of a $q$-type of a $k$-tuple $\bar{v}$ of vertices of a (colored) graph $G$, that is, the set $tp_{q}(G,\bar{u})$ of all first-order formulas $\xi(x_{1},\ldots,x_{k})$ of quantifier rank at most $q$ satisfying $G\models\xi(\bar{v})$. We note that the number of distinct $q$-type of $k$-tuple over a fixed signature is bounded by a function of $q$ and $k$. We refer the reader to Chapter I.3 of Achim Blumensath’s open access book [achim-book] for more details about finiteness of types.
We often need to work with types of sets rather that tuples. To that end, we say that a $k$-tuple $\bar{u}=(u_{1},\ldots,u_{k})$ is $\preceq$-ordered, if $u_{1}\preceq u_{2}\preceq\ldots\preceq u_{k}$ where $\preceq$ is a linear order. Observe that $\preceq$-ordered tuple $\bar{u}$ is uniquely determined by its set of vertices $\{u_{1},\ldots,u_{k}\}$ and their multiplicity.
Consider tuples of vertices of a graph $G$. We say that a tuple $\bar{u}$ $d$-separates tuples $\bar{v}_{1},\ldots,\bar{v}_{k}$ if, for every pair $i\neq j\in[k]$, every path $P$ of length at most $d$ between a vertex of $\bar{v}_{i}$ and a vertex of $\bar{v}_{j}$ contains a vertex of $\bar{u}$. Dreier [DBLP:conf/lics/Dreier21] proved a useful strengthening of Feferman-Vaught theorem showing how to recover the $q$-type of a tuple from $g(q)$-types of its subtuples:
Theorem 2.6 (Dreier [DBLP:conf/lics/Dreier21]).
There exists a function $g(q,\ell)$ such that for all labeled graphs $G$, every $q,\ell\in\mathbb{N}$, and all tuples $\bar{u},\bar{v}_{1},\ldots,\bar{v}_{k}$ of vertices from $G$ such that $\bar{u}$ $4^{q}$-separates $\bar{v}_{1},\ldots,\bar{v}_{k}$ and $|\bar{u}|+|\bar{v}_{1}|+\ldots+|\bar{v}_{k}|\leq\ell$, the type $tp_{q}(G,\bar{u}\bar{v}_{1}\ldots\bar{v}_{k})$ depends only on the types $tp_{g(q,\ell)}(G,\bar{u}\bar{v}_{1}),$ $\ldots,tp_{g(q,\ell)}(G,\bar{u}\bar{v}_{k})$. Furthermore, $tp_{q}(G,\bar{u}\bar{v}_{1}\ldots\bar{v}_{k})$ can be computed from $tp_{g(q,\ell)}(G,\bar{u}\bar{v}_{1}),\ldots,$ $tp_{g(q,\ell)}(G,\bar{u}\bar{v}_{k})$.
3 Transductions of classes admitting product structure
The following technical lemma constitutes the core of the proofs of our main results. We formulate it for 2-colored graphs (e.g., having black and white vertices) with the further intention to smoothly handle induced subgraphs.
Lemma 3.1.
There is a function $f:\mathbb{N}^{4}\to\mathbb{N}$ such that the following holds:
Let $\xi$ be a strongly $r$-local first-order binary formula of quantifier rank $q$ over the signature of 2-colored graphs. Let ${\cal Q}$ be a class of graphs of degree at most $d$. For any $Q\in{\cal Q}$ let $G$ be a 2-colored graph, such that $V(G)=V(Q\boxtimes M)$ and $E(G)\subseteq E(Q\boxtimes M)$, where $M$ has tree-width at most $k$.
Then, $\xi(G)$ has $\{Q^{\circ}_{r}\}$-clique-width at most $f(r,q,d,k)$, where $Q^{\circ}_{r}$ denotes the reflexive closure of the $r$-th power of $Q$.
Proof 3.2.
We postpone the definition of $f$ until the end of this proof. In the coming proof we will only consider the case that $r\geq 4^{q}$. This assumption is possible thanks to the following simple claim:
Claim 1.
Let $\phi$ be a $(Q^{\circ}_{\max\{r,4^{q}\}},f(r,q,d,k))$-expression valued $\xi(G)$. Then, the value of $\phi$ as a $(Q^{\circ}_{r},f(r,q,d,k))$-expression equals $\xi(G)$.
Proof 3.3 (Subproof).
We denote by $G_{r}$ the value of $\phi$ as a $(Q^{\circ}_{r},f(r,q,d,k))$-expression; this is sound since $Q^{\circ}_{\max\{r,4^{q}\}}$ and $Q^{\circ}_{r}$ have equal vertex set. Since $Q^{\circ}_{r}$ is a subgraph of $Q^{\circ}_{\max\{r,4^{q}\}}$, we get that $G_{r}$ is a subgraph of $\xi(G)$. Suppose for a contradiction that there is a pair of vertices $u=[p_{u},h_{u}],v=[p_{v},h_{v}]\in V(G_{r})=V(\xi(G))=V(G)$ such that $uv\in E(\xi(G))\setminus E(G_{r})$. Then, the distance in $Q$ between $p_{u}$ and $p_{v}$ is strictly more than $r$ (and at most $\max\{r,4^{q}\}$). Hence, the distance between $u$ and $v$ in $G$ is strictly greater than $r$, too, and so the edge $uv$ contradicts the assumption of strong $r$-locality of $\xi$.
Let ${\cal T}=(T,{\cal X}=(X_{t})_{t\in V(T)})$ be a smooth tree decomposition of $M$ of width $k$ rooted in $r_{0}\in V(T)$ such that the root bag $X_{r_{0}}=\emptyset$ is empty, every node of $T$ has at most two descendants, and for each leaf $t\in V(T)\setminus\{r_{0}\}$ the bag $X_{t}$ contains exactly one vertex. We observe that we can modify any tree decomposition to satisfy our requirements.
We denote by $t^{\uparrow}$ the parent node of $t$ in $T$. It follows from $X_{r_{0}}=\emptyset$ and from the interpolation property that for every vertex $h\in V(M)$ there is a unique node $t_{h}\in V(T)$ such that $h\in X_{t_{h}}\setminus X_{t_{h}^{\uparrow}}$. We say that $t_{h}$ is the forget node of $h$. Let $\preceq_{T}$ denote the leaves-to-root preorder of $T$. That is, $r_{0}\preceq_{T}t$ for all $t\in V(T)$. We fix an arbitrary linear order $\preceq_{M}$ satisfying that, for all $h_{1},h_{2}\in V(M)$, if $t_{h_{1}}\preceq_{T}t_{h_{2}}$ then $h_{1}\preceq_{M}h_{2}$.
We proceed in a way similar to the proof [DBLP:conf/mfcs/HlinenyJ24, Theorem 5.2]. However, roughly, instead of remembering adjacency to the not-yet-forgotten neighbors, we remember the $q^{\prime}$-type w.r. to the not-yet-forgotten vertices of the weak reachability sets for some constant $q^{\prime}$ depending only on $r$, $q$, $d$, and $k$.
For a node $t\in V(T)$, let $X^{+}_{t}\subseteq V(M)$ denote the union of $X_{s}$ where $s$ ranges over $t$ and all descendants of $t$. Let $Y_{t}=X^{+}_{t}\setminus X_{t^{\uparrow}}$ denote the vertices of $M$ which occur only in the bags of $t$ and its descendants. For the root $r_{0}$ of $T$, let specially $Y_{r_{0}}=X^{+}_{r_{0}}=V(M)$. Observe that all neighbors of a vertex $h\in Y_{t}$ in $V(M)\setminus Y_{t}$ must belong to the set $X_{t}\setminus Y_{t}$, by the interpolation property of a tree decomposition.
Claim 2.
Any path $P\subseteq M$ of length at most $r$ between vertices $h_{1}\in Y_{t}$ and $h_{2}\in V(M)\setminus Y_{t}$ intersects the set $\big{(}{\text{WReach}_{r}[M,\preceq_{M},h_{1}]}\cap X_{t}\big{)}\setminus Y_{t}$.
Proof 3.4 (Subproof).
It follows from the interpolation property and the previous observation that there is a vertex $h\in(X_{t}\setminus Y_{t})\cap V(P)$. Let $h$ be such vertex on the path $P$ closest to $h_{1}$, and let $P^{\prime}$ be the subpath of $P$ between $h_{1}$ and $h$. By the choices of $h$, we have $V(P^{\prime})\setminus\{h\}\subseteq Y_{t}$, and so $h\preceq_{M}h^{\prime}$ for all vertices $h^{\prime}\in V(P^{\prime})$. Hence, $h\in{\text{WReach}_{r}[M,\preceq_{M},h_{1}]}$.
Grohe et al. [DBLP:conf/wg/GroheKRSS15] proved that the weak $r$-coloring number of any graph $J$ is at most $wcol_{r}(J)\leq{r+tw(J)\choose tw(J)}$. Notably, our $\preceq_{M}$ is the order used in their proof to witness that $M$ has its weak $r$-coloring number at most (${r+k\choose k}$. That is, for every $h\in V(M)$, we get that $\big{|}{\text{WReach}_{r}[M,\preceq_{M},h]}\big{|}\leq{r+k\choose k}$.
We are approaching the core of the proof of Lemma 3.1. For an integer $\ell\geq 1$, we denote by $Q_{\ell}$ the $\ell$-th power of $Q$, and by $Q^{\circ}_{\ell}$ the reflexive closure of $Q_{\ell}$. We fix a proper coloring $s_{Q}:V(Q)\to[d^{3r}+1]$ of $Q_{3r}$. Such coloring exists since $Q$ has degree at most $d$, and so $Q_{3r}$ has degree at most $d^{3r}$.
Invariant. We are going to build a $(Q^{\circ}_{r},f(r,q,d,k))$-expression (cf. Definition 2.3) $\phi_{t}$ for each node $t\in V(T)$ using structural induction on $T$; maintaining the properties that
  1. (i)
    the value of $\phi_{t}$ is the subgraph $\xi(G)[V(Q)\times Y_{t}]$ of $\xi(G)$ induced by $V(Q)\times Y_{t}$, where
  2. (ii)
    the parameter vertex of each vertex $v=[p,h]\in V(Q)\times Y_{t}$ is $p\in V(Q^{\circ}_{r})=V(Q)$, and that
  3. (iii)
    the color $\hat{c}_{v,t}$ of $v$ is as prescribed by Definition 3.5.
Before giving the definition of the color $\hat{c}_{v,t}$, we recall the statement of Theorem 2.6, and set
$g^{\max{}}:=\max\{q,r,\max_{0\leq\,i\,\leq 2\left(d^{3r}+1\right)\cdot{r+k\choose k}}g(q,i)\},$ (1)
where $g$ is the function from Theorem 2.6.
Definition 3.5 (color $\hat{c}_{v,t}$).
Fix an arbitrary linear order $\preceq_{Q}$ on $V(Q)$ satisfying that, for all $p_{1},p_{2}\in V(Q)$, $s_{Q}(p_{1})\leq s_{Q}(p_{2})$ implies $p_{1}\preceq_{Q}p_{2}$. Let $\preceq_{G}$ be the linear order on $V(G)=V(Q)\times V(M)$ obtained by ordering the vertices lexicographically according to $\preceq_{M}$ and $\preceq_{Q}$, that is, $[p_{1},h_{1}]\preceq_{G}[p_{2},h_{2}]$ if and only if $(h_{1}\preceq_{M}h_{2}\land h_{1}\neq h_{2})\lor(p_{1}\preceq_{Q}p_{2}\land h_{1}=h_{2})$. Denote by $G^{s}$ the $2\!\left(d^{3r}+1\right)$-colored graph obtained from $G$ by coloring each vertex $v=[p,h]$ by the color $(c^{G}_{v},s_{Q}(p))$, where $c^{G}_{v}$ is the given color of $v$ in 2-colored $G$.
We define the initial color $\hat{c}_{v}$ of a vertex $v=[p_{v},h_{v}]\in V(Q)\times V(M)$ as a partial(!) function
$\hat{c}_{v}:2^{[d^{3r}+1]\times{\text{WReach}_{r}[M,\preceq_{M},h_{v}]}}\to\mbox{\it Types},$ (2)
where $\hat{c}_{v}(W)$ for $W\subseteq[d^{3r}+1]\times{\text{WReach}_{r}[M,\preceq_{M},h_{v}]}$ is partially defined as follows:
  • Let $Z_{v,W}:=\big{\{}u\!=\![p_{u},h_{u}]\in V(G)|\>p_{u}\in V(B_{r}(Q,p_{v}))$ $\wedge(s_{Q}(p_{u}),h_{u})\in W\big{\}}$. Let $\bar{w}_{v,W}$ be the unique $\preceq_{G}$-ordered tuple containing each vertex of $Z_{v,W}$ exactly once, and let $\bar{w}^{+}_{v,W}:=(\bar{w}_{v,W},v)$ denote this tuple after appending $v$ itself at the end. If $\{(s_{Q}(p),h)|$ $[p,h]\in Z_{v,W}\}=W$, [^3] As a brief explanation; each member of $W$ determines (in the set $Z_{v,W}$) at most one vertex of $G$ at close distance to $v$ (since $s_{Q}$ is a proper coloring of $Q_{3r}$). However, such vertex $u$ as determined by $(s_{Q}(p_{u}),h_{u})\in W$ may not exist in the graph $G$. In order to properly capture the relevant vertices $Z_{v,W}$ of $G$ with respect to $W$, we define $\hat{c}_{v}(W)$ only when the “completeness” condition $\{(s_{Q}(p),h)|$ $[p,h]\in Z_{v,W}\}=W$ is satisfied. then we let $\hat{c}_{v}(W)$ be the $g^{\max{}}$-type of $\bar{w}^{+}_{v,W}$ in $G^{s}$, and otherwise we leave $\hat{c}_{v}(W)$ undefined.
Subsequently, we define the running color $\hat{c}_{v,t}$ of a vertex $v=[p_{v},h_{v}]\in V(Q)\times V(M)$ at a node $t\in V(T)$ as the restriction of the initial color $\hat{c}_{v}$ of $v$ (above) to the sets ${W\subseteq\left[d^{3r}+1\right]\times\big{(}{\text{WReach}_{r}[M,\preceq_{M},h_{v}]}\setminus Y_{t}\big{)}}$.
Having defined the colors $\hat{c}_{c,t}$ (to be maintained in our construction), we will gradually describe the construction of the sought expressions $\phi_{t}$ on the vertex set $V(Q)\times Y_{t}$ for $t\in V(T)$. We begin by independently constructing their subexpressions $\psi_{h}$ valued the subgraph $\xi(G)[V(Q)\times\{h\}]$ for each vertex $h\in V(M)$ in 3. Figure 2 can be used as a “lightweight” (very simplified) illustration of this construction – steps (a)–(e) in Figure 2 illustrate building the expressions $\psi_{h}$, while steps (f)–(l) in the picture correspond to the subsequent construction of the expressions $\phi_{t}$ below.
Claim 3.
There is a function $f^{\prime}:\mathbb{N}^{4}\to\mathbb{N}$ such that, for each vertex $h\in V(M)$, there is a $(Q^{\circ}_{r},f^{\prime}(r,q,d,k))$-expression $\psi_{h}$ valued $\xi(G)[V(Q)\times\{h\}]$, such that each vertex $v=[p_{v},h]\in V(Q)\times\{h\}$ has its initial color $\hat{c}_{v}$ and parameter vertex $p_{v}$.
Proof 3.6 (Subproof).
For each nonempty subset $Z\subseteq V(Q)$, by induction on $|Z|$, we build a $(Q^{\circ}_{r},f(r,q,d,k))$-expression $\psi_{h}^{Z}$ valued $\xi(G)[Z\times\{h\}]$, such that each vertex $v=[p_{v},h]\in Z\times\{h\}$ has color $(s_{Q}(p_{v}),\hat{c}_{v})$ and parameter vertex $p_{v}$.
In the base case, that is $|Z|=1$, the expression $\psi_{h}^{Z}$ is trivial since we only create a single labeled vertex without edges.
Assume $|Z|\geq 2$ and choose a vertex $v=[p_{v},h]$ where $p_{v}\in Z$. We apply our induction assumption to obtain the expression $\psi_{h}^{Z\setminus\{p_{v}\}}$ and denote by $K$ its value. We make $K+v$ by taking a disjoint union of $\psi_{h}^{Z\setminus\{p_{v}\}}$ and an expression creating $v$ with parameter vertex $p_{v}$ and a special color NewVertex. Then, for every $u=[p_{u},h]\in({Z\setminus\{p_{v}\}})\times\{h\}$ such that $uv\in E(\xi(G))$, we apply the operation of adding edges (cf. Definition 2.3.d) between colors NewVertex and $(s_{Q}(p_{u}),\hat{c}_{u})$.
A crucial observation is that, by doing so, we create only one edge for each such choice of $u$, that is, the edge $uv$. Indeed, first note that since $\xi$ is strongly $r$-local, the distance between $u$ and $v$ in $G$, as well as the distance between $p_{u}$ and $p_{v}$ in $Q$, is at most $r$, and so $p_{u}p_{v}\in E(Q_{r}^{\circ})$. Let $x=[p_{x},h]$ be any vertex of color equal to $(s_{Q}(p_{u}),\hat{c}_{u})$ whose parameter vertex $p_{x}$ is adjacent to $p_{v}$ in $Q^{\circ}_{r}$. Then the distance in $Q$ between $p_{x}$ and $p_{v}$ is also at most $r$, and hence between $p_{x}$ and $p_{u}$ at most $2r$ by the triangle inequality. Since $s_{Q}$ is a proper coloring of $Q_{3r}$ and $s_{Q}(p_{x})=s_{Q}(p_{u})$, we conclude $p_{x}=p_{u}$, and consequently, $x=u$.
Therefore, we have created an expression valued $\xi(G)[Z\times\{h\}]$, and now we recolor NewVertex to $(s_{Q}(p_{v}),\hat{c}_{v})$ to finish the desired expression $\psi_{h}^{Z}$. Finally, we obtain $\psi_{h}$ from $\psi_{h}^{V(Q)}$ by recoloring each vertex from any color $(a,c)$ to color $c$.
As the last task, we confirm that the number of colors used in the expression $\psi_{h}$ is bounded by a function ($f^{\prime}$) of $r$, $q$, $d$, and $k$, because the size of the codomain of $s_{Q}$ is at most $d^{3r}+1$, and unions of both domains and codomains of the functions $\hat{c}_{u}$ over all vertices $u\in V(Q)\times\{h\}$ have bounded size. Indeed, all the domains are subsets of $\left[d^{3r}+1\right]\times{\text{WReach}_{r}[M,\preceq_{M},h]}$, and so their union is of bounded size. The number of possible types in the codomains of functions $\hat{c}_{u}$ is bounded, since the number of colors is bounded, the quantifier rank is bounded, and the sizes of the tuples are bounded.
We now build the sought expression $\phi_{t}$. Assume that, for each child $t^{\prime}$ of $t$, we have inductively built expressions $\phi_{t^{\prime}}$ satisfying the aforementioned invariants (i), (ii), and (iii). We also shortly denote by $V(\phi)$ the vertex set of the value of $\phi$. Then, we proceed with the following steps:
  1. 1.
    If $t$ is a leaf, then $X_{t}=\{h\}$ for some node $h\in V(M)$. So, we create $\phi_{t}$ from $\psi_{h}$ by recoloring each vertex $w$ from its initial color $\hat{c}_{w}$ to its running color $\hat{c}_{w,t}$.
  2. 2.
    We further assume that $t$ is not a leaf, and denote by $t_{i}$ for $i=1$ or $i\in\{1,2\}$ the children of $t$ in $T$. Let $\phi_{t_{i}}^{\uparrow}$ be the expression obtained from $\phi_{t_{i}}$ by recoloring each color $c$ to $(i,c)$ in $\phi_{t_{i}}^{\uparrow}$. If $t$ has only one child, we set $\phi_{t}^{1}:=\phi_{t_{1}}^{\uparrow}$. Otherwise, $t$ has exactly two children $t_{1}$ and $t_{2}$, and we create $\phi_{t}^{1}$ by taking the disjoint union of $\phi_{t_{1}}^{\uparrow}$ and $\phi_{t_{2}}^{\uparrow}$.
  3. 3.
    If $Y_{t}\setminus(Y_{t_{1}}\cup Y_{t_{2}})=\emptyset$, then we set $\phi_{t}^{2}:=\phi_{t}^{1}$. Otherwise, it follows from smoothness of the decomposition of $M$ that $Y_{t}\setminus(Y_{t_{1}}\cup Y_{t_{2}})=\{h\}$ for some $h\in V(M)$. We create an expression $\psi_{h}^{\uparrow}$ from $\psi_{h}$ by recoloring each color $c$ to $(3,c)$, and $\phi_{t}^{2}$ by taking the disjoint union of $\phi_{t}^{1}$ and $\psi_{h}^{\uparrow}$.
  4. 4.
    We create $\phi_{t}^{3}$ from $\phi_{t}^{2}$ by adding edges between the following pairs of colors:
    1. (A)
      If $t$ has two children, then we do the following: For every vertex pair $u\in V(Q)\times Y_{t_{1}}=V(\phi_{t_{1}})$ and $v\in V(Q)\times Y_{t_{2}}=V(\phi_{t_{2}})$ such that $uv\in E(\xi(G))$, we apply the operation of adding edges between colors $(1,\hat{c}_{u,t_{1}})$ and $(2,\hat{c}_{v,t_{2}})$ (cf. Definition 2.3.d).
    2. (B)
      If $Y_{t}\setminus(Y_{t_{1}}\cup Y_{t_{2}})\neq\emptyset$, that is, $Y_{t}\setminus(Y_{t_{1}}\cup Y_{t_{2}})=\{h\}$, then we do the following: For every $i\in\{1,2\}$ and every pair $u\in V(\phi_{t_{i}})$ and $w\in V(Q)\times\{h\}=V(\psi_{h})$ such that $uw\in E(\xi(G))$, we apply the operation of adding edges between colors $(i,\hat{c}_{u,t_{i}})$ and $(3,\hat{c}_{w})$.
  5. 5.
    Finally, we make $\phi_{t}$ from $\phi_{t}^{3}$ by performing the following recoloring operations to satisfy invariant (iii):
    • For $i=1,2$, replace each color $(i,\hat{c}_{u,t_{i}})$ by $\hat{c}_{u,t}$.
    • Replace each color $(3,\hat{c}_{w})$ by $\hat{c}_{w,t}$.
We note that the recoloring operations of the steps 1. and 5. are well-defined, because the way we restrict the function $\hat{c}_{u,t_{i}}$ to obtain $\hat{c}_{u,t}$ (resp. $\hat{c}_{w}$ to obtain $\hat{c}_{w,t}$) depends only on $t$.
We denote by $G^{t}$ the value of $\phi_{t}$. It is clear that $V(G^{t})=V(Q)\times Y_{t}$, and that the colors and the parameter vertices of $\phi_{t}$ satisfy our invariants (ii) and (iii). We continue with:
Claim 4.
$E\big{(}\xi(G)[V(Q)\times Y_{t}]\big{)}\subseteq E(G^{t})$
Proof 3.7 (Subproof).
Suppose otherwise, and let $uv\in E(\xi(G)[V(Q)\times Y_{t}])\setminus E(G^{t})$ be an edge we have not created. If $u,v\in V(\phi_{t_{1}})$ or $u,v\in V(\phi_{t_{2}})$, then we obtain a contradiction with the induction assumption. Similarly, if $u,v\in V(\psi_{h})$, then we contradict 3. Hence, $uv$ is between two of the three subgraphs in the previous construction. It follows from strong $r$-locality of $\xi$ that the distance between $u$ and $v$ in $G$ is at most $r$, and so is the distance between their parameter vertices in $Q$, and the parameter vertices are adjacent in $Q^{\circ}_{r}$. Hence, we have created the edge $uv$ in step 4. above.
Claim 5.
$E(G^{t})\subseteq E\big{(}\xi(G)[V(Q)\times Y_{t}]\big{)}$
Proof 3.8 (Subproof).
Suppose otherwise for a contradiction. Let $xy\in E(G^{t})\setminus E(\xi(G)[V(Q)\times Y_{t}])$ be any unwanted edge of $G^{t}$. As in 4, we must have added the edge $xy$ in step 4.. Then, there is an edge $uv\in E(\xi(G)[V(Q)\times Y_{t}])$ such that $x$ and $u$ as well as $y$ and $v$ have the same color. Let $[p_{x},h_{x}]=x$, $[p_{y},h_{y}]=y$, $[p_{u},h_{u}]=u$, and $[p_{v},h_{v}]=v$.
It follows from strong $r$-locality of $\xi$ that the distance between $u$ and $v$ in $G$ it at most $r$, and so the distance between $h_{u}$ and $h_{v}$ in $M$ is at most $r$. Furthermore, we have created both edges $xy$ and $uv$, and so $p_{u}p_{v}\in E(Q_{r}^{\circ})$ and $p_{x}p_{y}\in E(Q_{r}^{\circ})$ and, consequently, the $Q$-distance within each of the pairs $p_{u},p_{v}$ and $p_{x},p_{y}$ is at most $r$.
[…]
  • (I)
    $W_{a,b}$ is the unique inclusion-maximal subset of $[d^{3r}+1]\times V(M)$ such that both values $\tilde{c}_{a}(W_{a,b})$ and $\tilde{c}_{b}(W_{a,b})$ are defined.
  • (II)
    Both $\preceq_{G}$-ordered tuples $\bar{w}_{a,W_{a,b}}$ and $\bar{w}_{b,W_{a,b}}$ contain the same set of vertices $Z^{\prime}_{a,b}$.
  • (III)
    The set $Z^{\prime}_{a,b}$ is an $r$-separator of $a$ and $b$.
We distinguish two cases, depending on whenever the edge $xy$ was created in step 4.4.(A) or step 4.4.(B).
Case 1.
The edge $xy$ has been added in step 4.4.(B).
Without loss of generality, we assume that $x,u\in V(\phi_{t_{i}})$ where $i\in\{1,2\}$, and $y,v\in V(\psi_{h})$ where $h=h_{y}=h_{v}$. Then $\tilde{c}_{x}=\hat{c}_{x,t_{i}}$, $\tilde{c}_{y}=\hat{c}_{y}$, $\tilde{c}_{u}=\hat{c}_{u,t_{i}}$, and $\tilde{c}_{v}=\hat{c}_{v}$.
Recalling Definition 3.5, we observe that $x$ and $u$ (resp., $y$ and $v$) have the same type in the colored graph $G^{s}$, because $\hat{c}_{x,t_{i}}(\emptyset)=\hat{c}_{u,t_{i}}(\emptyset)$ (resp., $\hat{c}_{y}(\emptyset)=\hat{c}_{v}(\emptyset)$) is always defined, and the value of $\hat{c}_{z}(\emptyset)=\hat{c}_{z,t^{\prime}}(\emptyset)$ is the type of $z$ for all vertices $z\in V(G)$ and all nodes $t^{\prime}\in V(T)$. In particular, since each vertex color in $G^{s}$ contains the color of its parameter vertex, we conclude that $s_{Q}(p_{x})=s_{Q}(p_{u})$ and $s_{Q}(p_{y})=s_{Q}(p_{v})$.
We continue the proof with three technical subclaims. For any $z=[p_{z},h_{z}]\in V(\phi_{t_{i}})$, let us shortly denote $Z^{1}_{z}:=V(B_{r}(Q,p_{z}))\times\big{(}{\text{WReach}_{r}[M,\preceq_{M},h_{z}]}\setminus Y_{t_{i}}\big{)}$. Similarly, for any $z=[p_{z},h_{z}]\in V(\psi_{h})$, we shortly denote $Z^{2}_{z}:=V(B_{r}(Q,p_{z}))\times{\text{WReach}_{r}[M,\preceq_{M},h_{z}]}$.
Given any vertices $a=[p_{a},h_{a}]\in V(\phi_{t_{i}})$ and $b=[p_{b},h_{b}]\in V(\psi_{h})$, we denote by $Z^{\prime}_{a,b}$ the set $Z^{\prime}_{a,b}:=Z^{1}_{a}\cap Z^{2}_{b}$. Let $W_{a,b}:=\{(s_{Q}(p),h)|\>(p,h)\in Z^{\prime}_{a,b}\}$, and let $D_{a,b}\subseteq\mbox{$[d^{3r}+1]$}\times V(M)$ be an arbitrary inclusion-maximal set such that both values $\hat{c}_{a,t_{i}}(D_{a,b})$ and $\hat{c}_{b}(D_{a,b})$ are defined.
In this case, conditions (I), (II), and (III) are fulfilled in order by the following three claims:
Claim 6.
Assume that distance between $p_{a}$ and $p_{b}$ is at most $r$ in $Q$. Then $D_{a,b}=W_{a,b}$, and so $D_{a,b}$ is uniquely determined.
Proof 3.9 (Subproof).
Let $(c,h^{\prime})\in D_{a,b}$ be arbitrary. Then, the set $Z_{a,D_{a,b}}$ (cf. Definition 3.5) contains a vertex $[p^{a},h^{\prime}]$ such that $c=s_{Q}(p^{a})$ and $p^{a}\in V(B_{r}(Q,p_{a}))$. Furthermore, $h^{\prime}\in{\text{WReach}_{r}[M,\preceq_{M},h_{a}]}\setminus Y_{t_{i}}$, because $\hat{c}_{a,t_{i}}$ is the restriction of $\hat{c}_{a}$ to subsets of $[d^{3r}+1]\times({\text{WReach}_{r}[M,\preceq_{M},h_{a}]}\setminus Y_{t_{i}})$. So, $[p^{a},h^{\prime}]\in Z^{1}_{a}$.
Similarly, the set $Z_{b,D_{a,b}}$ contains a vertex $[p^{b},h^{\prime}]$ such that $c=s_{Q}(p^{b})$ and $p^{b}\in V(B_{r}(Q,p_{b}))$. Furthermore, $h^{\prime}\in{\text{WReach}_{r}[M,\preceq_{M},h_{b}]}$ since the domain of $\hat{c}_{b}$ contains only subsets of $[d^{3r}+1]\times{\text{WReach}_{r}[M,\preceq_{M},h_{b}]}$. So, $[p^{b},h^{\prime}]\in Z^{2}_{b}$.
The distance in $Q$ between $p_{a}$ and $p_{b}$ is at most $r$, and since $s_{Q}$ is a proper coloring of $Q_{3r}$, we get that $p^{a}=p^{b}$. Hence, $[p^{a},h]=[p^{b},h]\in Z^{\prime}_{a,b}$, so $(c,h)=(s_{Q}(p^{a}),h)\in W_{a,b}$.
Altogether, $D_{a,b}\subseteq W_{a,b}$. We observe that both $\hat{c}_{a}(W_{a,b})$ and $\hat{c}_{b}(W_{a,b})$ are defined as witnessed by the vertices in $Z^{\prime}_{a,b}$. Hence, $W_{a,b}\subseteq D_{a,b}$.
Claim 7.
Both tuples $\bar{w}_{a,W_{a,b}}$ and $\bar{w}_{b,W_{a,b}}$ contain exactly the vertices of $Z^{\prime}_{a,b}$.
Proof 3.10 (Subproof).
This follows from Definition 3.5 because for any two vertices $c=[p_{c},h_{c}]$ and $d=[p_{d},h_{d}]$ satisfying $(s_{Q}(p_{c}),h_{c})=(s_{Q}(p_{d}),h_{d})$ and $p_{c},p_{d}\in V(B_{r}(Q,p_{a}))$, we directly get $h_{c}=h_{d}$, and $p_{c}=p_{d}$ (so, $c=d$) since $s_{Q}$ is a proper coloring of $Q_{3r}$ and $p_{c}p_{d}\in E(Q_{3r})$ if not equal.
Claim 8.
The set $Z^{\prime}_{a,b}$ $r$-separates $a$ and $b$ in $G$.
Proof 3.11 (Subproof).
Recall that $a\in V(\phi_{t_{i}})$ and $b\in V(\psi_{h})$. We assume that the distance between $a$ and $b$ in $G$ is at most $r$; otherwise, the claim is trivial. Consider any path $P\subseteq G$ of length at most $r$ between $a$ and $b$. Recall that the mapping $[q,h]\mapsto h$ from vertices of $G$ to their parameter vertices in the expression is a graph homomorphism, that is, it preserves the edge relation. We denote by $P_{M}$ the image of $P$ in $M$ under this homomorphism.
Let $h_{0}$ be the $\preceq_{M}$-minimal vertex of $P_{M}$. Then, $P_{M}$ certifies that $h_{0}\in{\text{WReach}_{r}[M,\preceq_{M},h_{a}]}\cap{\text{WReach}_{r}[M,\preceq_{M},h_{b}]}$. It follows from the interpolation property that $P_{M}$ contains a vertex $h^{\prime}\in X_{t}$. Notably, $h^{\prime}\not\in Y_{t_{i}}$ and $h_{0}\preceq_{M}h^{\prime}$, so $h_{0}\not\in Y_{t_{i}}$. Let $z_{0}=[p_{0},h_{0}]$ be an arbitrary preimage of $h_{0}$ in $G$. Furthermore, since the distance in $G$ between each of $a$, $b$ and $z_{0}$ is at most $r$, we get that the distance between each of $p_{a}$, $p_{b}$ and $p_{0}$ in $Q$ is at most $r$, and so $p_{0}\in V(B_{r}(Q,p_{a}))\cap V(B_{r}(Q,p_{b}))$. Therefore, by the definition, $z_{0}\in Z^{\prime}_{a,b}$.
Case 2.
The edge $xy$ has been added in step 4.4.(A).
Without loss of generality, we now assume that $x,u\in V(\phi_{t_{1}})$ and $y,v\in V(\phi_{t_{2}})$. Then $\tilde{c}_{x}=\hat{c}_{x,t_{1}}$, $\tilde{c}_{y}=\hat{c}_{y,t_{2}}$, $\tilde{c}_{u}=\hat{c}_{u,t_{1}}$, and $\tilde{c}_{v}=\hat{c}_{v,t_{2}}$.
We observe, as in 1 again, that $x$ and $u$ (resp. $y$ and $v$) have the same type in $G^{s}$, because $\hat{c}_{x,t_{1}}(\emptyset)=\hat{c}_{u,t_{1}}(\emptyset)$ (resp. $\hat{c}_{y,t_{2}}(\emptyset)=\hat{c}_{v,t_{2}}(\emptyset)$) is always defined, and the value of $\hat{c}_{z,t^{\prime}}(\emptyset)$ is the type of $z$ for all vertices $z\in V(G)$ and all nodes $t^{\prime}\in V(T)$. So, again, $s_{Q}(p_{x})=s_{Q}(p_{u})$ and $s_{Q}(p_{y})=s_{Q}(p_{v})$.
We again continue the proof with three technical subclaims corresponding in order to conditions (I), (II), and (III), as stated before 1.
As in 1, this time for both $i\in\{1,2\}$ and any $z=[p_{z},h_{z}]\in V(\phi_{t_{i}})$, we shortly write $Z^{i}_{z}:=V(B_{r}(Q,p_{z}))\times\big{(}{\text{WReach}_{r}[M,\preceq_{M},h_{z}]}\setminus Y_{t_{i}}\big{)}$. Given any vertices $a=[p_{a},h_{a}]\in V(\phi_{t_{1}})$ and $b=[p_{b},h_{b}]\in V(\phi_{t_{2}})$, we denote by $Z^{\prime}_{a,b}$ the set $Z^{\prime}_{a,b}:=Z^{1}_{a}\cap Z^{2}_{b}$. Let $W_{a,b}:=\{(s_{Q}(p),h)|\>(p,h)\in Z^{\prime}_{a,b}\}$, and let $D_{a,b}\subseteq\mbox{$[d^{3r}+1]$}\times V(M)$ be an arbitrary inclusion-maximal set such that both values $\hat{c}_{a,t_{1}}(D_{a,b})$ and $\hat{c}_{b,t_{2}}(D_{a,b})$ are defined.
Claim 9.
Assume that distance between $p_{a}$ and $p_{b}$ is at most $r$ in $Q$. Then $D_{a,b}=W_{a,b}$, and so $D_{a,b}$ is uniquely determined.
Proof 3.12 (Subproof).
Let $(c,h)\in D_{a,b}$ be arbitrary. Then, the set $Z_{a,D_{a,b}}$ (cf. Definition 3.5) contains a vertex $[p^{a},h]$ such that $c=s_{Q}(p^{a})$ and $p^{a}\in V(B_{r}(Q,p_{a}))$. Furthermore, $h\in{\text{WReach}_{r}[M,\preceq_{M},h_{a}]}\setminus Y_{t_{1}}$, because $\hat{c}_{a,t_{1}}$ is the restriction of $\hat{c}_{a}$ to subsets of $[d^{3r}+1]\times({\text{WReach}_{r}[M,\preceq_{M},h_{a}]}\setminus Y_{t_{1}})$. So, $[p^{a},h]\in Z^{1}_{a}$.
We analogously obtain that the set $Z_{b,D_{a,b}}$ contains a vertex $[p^{b},h]$ such that $c=s_{Q}(p^{b})$, $p^{b}\in V(B_{r}(Q,p_{b}))$, and $[p^{b},h]\in Z^{2}_{b}$. The distance in $Q$ between $p_{a}$ and $p_{b}$ is at most $r$, and since $s_{Q}$ is a proper coloring of $Q_{3r}$, we get that $p^{a}=p^{b}$. Hence, $[p^{a},h]=[p^{b},h]\in Z^{\prime}_{a,b}$, and so $(c,h)=(s_{Q}(p^{a}),h)\in W_{a,b}$.
Altogether, $D_{a,b}\subseteq W_{a,b}$. We observe that both $\hat{c}_{a}(W_{a,b})$ and $\hat{c}_{b}(W_{a,b})$ are defined as witnessed by the vertices in $Z^{\prime}_{a,b}$. Hence, $W_{a,b}\subseteq D_{a,b}$.
Claim 10.
Both tuples $\bar{w}_{a,W_{a,b}}$ and $\bar{w}_{b,W_{a,b}}$ contain exactly the vertices of $Z^{\prime}_{a,b}$.
Proof 3.13 (Subproof).
This follows from Definition 3.5 because for any two vertices $c=[p_{c},h_{c}]$ and $d=[p_{d},h_{d}]$ satisfying $(s_{Q}(p_{c}),h_{c})=(s_{Q}(p_{d}),h_{d})$ and $p_{c},p_{d}\in V(B_{r}(Q,p_{a}))$, we directly get $h_{c}=h_{d}$, and $p_{c}=p_{d}$ (so, $c=d$) since $s_{Q}$ is a proper coloring of $Q_{3r}$ and $p_{c}p_{d}\in E(Q_{3r})$ if not equal.
Claim 11.
The set $Z^{\prime}_{a,b}$ $r$-separates $a$ and $b$ in $G$.
Proof 3.14 (Subproof).
Recall that $a\in V(\phi_{t_{1}})$ and $b\in V(\phi_{t_{2}})$. We assume that the distance between $a$ and $b$ in $G$ is at most $r$; otherwise, the claim is trivial. Consider any path $P\subseteq G$ of length at most $r$ between $a$ and $b$. Recall that the mapping $[q,h]\mapsto h$ from vertices of $G$ to their parameter vertices in the expression is a graph homomorphism, that is, it preserves the edge relation. We denote by $P_{M}$ the image of $P$ in $M$ under this homomorphism.
Let $h_{0}$ be the $\preceq_{M}$-minimal vertex of $P_{M}$. Then, $P_{M}$ certifies that $h_{0}\in{\text{WReach}_{r}[M,\preceq_{M},h_{a}]}\cap{\text{WReach}_{r}[M,\preceq_{M},h_{b}]}$. It follows from the interpolation property that $P_{M}$ contains a vertex $h^{\prime}\in X_{t}$. Notably, $h^{\prime}\not\in Y_{t_{1}}\cup Y_{t_{2}}$ and $h_{0}\preceq_{M}h^{\prime}$, so $h_{0}\not\in Y_{t_{1}}\cup Y_{t_{2}}$. Let $z_{0}=[p_{0},h_{0}]$ be an arbitrary preimage of $h_{0}$ in $G$. Furthermore, since the distance in $G$ between each of $a$, $b$ and $z_{0}$ is at most $r$, we get that the distance between each of $p_{a}$, $p_{b}$ and $p_{0}$ in $Q$ is at most $r$, and so $p_{0}\in V(B_{r}(Q,p_{a}))\cap V(B_{r}(Q,p_{b}))$. Therefore, by the definition, $z_{0}\in Z^{\prime}_{a,b}$.
Back in 5, since $\tilde{c}_{u}(W_{u,v})$ and $\tilde{c}_{v}(W_{u,v})$ are defined, we then get that $\tilde{c}_{x}(W_{u,v})=\tilde{c}_{u}(W_{u,v})$ and $\tilde{c}_{y}(W_{u,v})=\tilde{c}_{v}(W_{u,v})$ are defined as well. Hence, using (I), we get that $W_{u,v}\subseteq W_{x,y}$. Similarly, since $\tilde{c}_{x}(W_{x,y})$ and $\tilde{c}_{y}(W_{x,y})$ are defined, we get that $\tilde{c}_{u}(W_{x,y})=\tilde{c}_{x}(W_{x,y})$ and $\tilde{c}_{v}(W_{x,y})=\tilde{c}_{y}(W_{x,y})$ are defined as well. Hence, $W_{x,y}\subseteq W_{u,v}$. Altogether, $W_{x,y}=W_{u,v}$.
Hence, $\bar{w}_{u,W_{u,v}}=\bar{w}_{v,W_{u,v}}$ (since they are the same $\preceq_{G}$-ordering of the same set $Z^{\prime}_{u,v}$ by (II)), and similarly $\bar{w}_{x,W_{u,v}}=\bar{w}_{x,W_{x,y}}=\bar{w}_{y,W_{x,y}}=\bar{w}_{y,W_{u,v}}$. Notably, the tuples $\bar{w}_{u,W_{u,v}}=\bar{w}_{v,W_{u,v}}$ and $\bar{w}_{x,W_{x,y}}=\bar{w}_{y,W_{x,y}}$ have the same size equal to $|W_{u,v}|=|W_{x,y}|$ (see Definition 3.5).
Recall that $r\geq 4^{q}$. From (II) and (III), we know that $\bar{w}_{u,W_{u,v}}=\bar{w}_{v,W_{u,v}}$ $4^{q}$-separates $u$ and $v$, and likewise $\bar{w}_{x,W_{x,y}}=\bar{w}_{y,W_{x,y}}$ $4^{q}$-separates $x$ and $y$. So, we apply Theorem 2.6 to obtain that the type $tp_{q}(G^{s},uv)$ of the pair $uv$ (resp. the type $tp_{q}(G^{s},xy)$ of $xy$) depends only on the types $tp_{g^{\max}}(G^{s},\bar{w}_{u,W_{uv}}u)=\tilde{c}_{u}(W_{uv})=\tilde{c}_{x}(W_{xy})=tp_{g^{\max}}(G^{s},\bar{w}_{x,W_{xy}}x)$ and $tp_{g^{\max}}(G^{s},\bar{w}_{v,W_{uv}}v)=\tilde{c}_{v}(W_{uv})$ $=\tilde{c}_{y}(W_{xy})=tp_{g^{\max}}(G^{s},\bar{w}_{x,W_{xy}}x)$; hence $tp_{q}(G^{s},uv)=tp_{q}(G^{s},xy)$. In particular, $\xi\in tp_{q}(G^{s},uv)=tp_{q}(G^{s},xy)$, and so $xy\in E(\xi(G))$, a contradiction.
This finishes the proof of 5.
We have shown that $\phi_{t}$ satisfies all three invariants. However, it remains to show that the number of colors is bounded by a function $f$ of $r$, $q$, $d$, and $k$. We remark that the whole set $\{\hat{c}_{v,t}|v\in V(G),t\in V(T)\}$ has unbounded size. We show that, for each subexpression $\phi$ of $\phi_{t}$, the value of $\phi$ contains only a bounded number of colors. By doing so, we implicitly define the function $f$. Once we have done so, we show how to replace the functions $\hat{c}_{v,t}$ by numbers from the set $[f(r,q,d,k)]$ so that we obtain an expression with the same value and bounded total number of colors.
Claim 12.
For each subexpression $\phi$ of $\phi_{t}$, the value of $\phi$ features a number of colors at most $f(r,q,d,k)$, where the function $f$ depending only on $r$, $q$, $d$, $k$ is defined in the proof.
Proof 3.15 (Subproof).
By the induction assumption, as expressed in 3, the subexpressions $\psi_{h}$, $\phi_{t_{1}}$, and $\phi_{t_{2}}$ contain at most $f(r,q,d,k)$ (so-far implicit number) of colors.
We now bound the number of colors involved when creating $\phi_{t}$ from $\psi_{h}$, $\phi_{t_{1}}$, and $\phi_{t_{2}}$. The worst case, over all possible values of these subexpressions, defines the function $f(r,q,d,k)$.
We observe that it suffices to bound the size of both unions of domains and codomains of functions $\hat{c}_{a}$, $\hat{c}_{a,t}$, $\hat{c}_{a,t_{1}}$, and $\hat{c}_{a,t_{2}}$. Size of the union of the codomains is clearly bounded, since there is only bounded number of $g^{\max}$-types of tuples of size at most $2\left(d^{3r}+1\right)\cdot{r+k\choose k}$. We have already shown that the union of the domains of $\hat{c}_{a,t}$ has bounded size in the proof of 3, and we skip repetition of the argument here. We observe that, in order to bound the size of unions of the domains of the remaining functions, it suffices to bound the size of the set $\bigcup_{h\in Y_{t}}{\text{WReach}_{r}[M,\preceq_{M},h]}\setminus Y_{t^{\prime}}$ for all $t^{\prime}\in\{t,t_{1},t_{2}\}$. Let $t^{\prime}\in\{t,t_{1},t_{2}\}$ be arbitrary. We show that $\bigcup_{h\in Y_{t^{\prime}}}{\text{WReach}_{r}[M,\preceq_{M},h]}\setminus Y_{t^{\prime}}\subseteq\bigcup_{h\in X_{t^{\prime}}}{\text{WReach}_{r}[M,\preceq_{M},h]}$. Let $h_{1}\in Y_{t^{\prime}}$ be arbitrary. Observe that, if there is a vertex $h_{2}\in{\text{WReach}_{r}[M,\preceq_{M},h]}\setminus Y_{t^{\prime}}$, then it follows from 2 that there is a vertex $p\in X_{t^{\prime}}$ such that $h_{2}\in{\text{WReach}_{r}[M,\preceq_{M},p]}$. Hence,
$\left|\bigcup_{h\in Y_{t^{\prime}}}{\text{WReach}_{r}[M,\preceq_{M},h]}\setminus Y_{t^{\prime}}\right|\leq$ $\left|\bigcup_{h\in X_{t^{\prime}}}{\text{WReach}_{r}[M,\preceq_{M},h]}\right|\leq(k+1)\cdot{r+k\choose k}$.
We have shown that each subexpression uses only a bounded number of colors. Our upper bounds depend only on $r$, $q$, $d$, and $k$, so we have properly defined the sought function $f$.
We show how to replace the colors $\hat{c}_{v,t}$ by numbers from the set $[f(r,q,d,k)]$. We do so recursively, top-down, starting at the root $r_{0}$ of $T$ and the expression $\phi_{r_{0}}$. Let $\phi$ be arbitrary subexpression of $\phi_{r_{0}}$ valued $K$. If $\phi=\phi_{r_{0}}$, then we choose an arbitrary injection from the colors used by $K$ to the set $[f(r,q,d,k)]$. It follows from 12 that such injective function exists. If $\phi$ has a parent subexpression $\phi^{\uparrow}$ valued $K^{\uparrow}$, then we assume that we have already defined an injective function $o^{\uparrow}$ mapping colors used by the vertices of $K^{\uparrow}$ to the numbers from the set $[f(r,q,d,k)]$. We define analogous function $o$ for $\phi$, such that $o$ and $o^{\uparrow}$ agree on all common colors. That is, let $c$ be arbitrary color used by at least one vertex of $K$. If $c$ is also used in $K^{\uparrow}$, then we set $o(c):=o^{\uparrow}(c)$. Otherwise, we have obtained $\phi^{\uparrow}$ from $\phi$ by recoloring $c$ to another color $c^{\uparrow}$. If $c^{\uparrow}$ is not used in the coloring of $K$, we set $o(c):=o^{\uparrow}(c^{\uparrow})$. Otherwise, $c^{\uparrow}$ is used in $K$, so $K^{\uparrow}$ uses at most $f(r,q,d,k)-1$ colors (that is, at most one color less than $K$). Since there can be only one such color $c$ (because Definition 2.3 allows only one recoloring at a time), we set $o(c)$ to the number from the set $[f(r,q,d,k)]$ unused by $o^{\uparrow}$.
Finally, in each subexpression, we replace the colors by the assigned numbers from the set $[f(r,q,d,k)]$ according to the above constructed assignment. By doing so, we obtain an expression with the same value (except for the color names).
Since $X_{r_{0}}=\emptyset$, we get that $Y_{r_{0}}=V(M)$, so $V(Q)\times Y_{r_{0}}=V(G)$, and it follows from the invariant (iii) that the value of $\phi_{r_{0}}$ is $\xi(G)$. Hence, we obtain a $(Q_{r}^{\circ},f(r,q,d,k))$-expression for $\xi(G)$, thus finishing the proof of Lemma 3.1.
Lastly, we derive the two main theorems of this article.
Theorem 3.16.
Assume a class ${\cal Q}$ of bounded-degree graphs, and let ${\cal C}$ be a class of graphs admitting ${\cal Q}$-product structure. Let $\tau$ be a first-order transduction such that $\tau({\cal C})$ is a class of simple graphs. Then, for some integer $r$ depending only on $\tau$, the following holds. Denoting by ${\cal Q}^{\circ}_{r}$ the class containing the reflexive closures of $r$-th powers of graphs from ${\cal Q}$, the class $\tau({\cal C})$ is a perturbation of a class of bounded ${\cal Q}^{\circ}_{r}$-clique-width.
Furthermore, if $\tau$ is immersive, then $\tau({\cal C})$ itself is of bounded ${\cal Q}^{\circ}_{r}$-clique-width.
Proof 3.17.
Recall Theorem 2.5 showing that every first-order transduction $\tau$ is subsumed by a composition of a copying operation $C$, an immersive transduction $\tau^{\prime}$ and a perturbation $P$. Since $\tau({\cal C})$ is a class of simple graphs, we get that $\tau^{\prime}$ contains only a single formula $\xi^{\prime}$. Furthermore, it follows from immersiveness that $\xi^{\prime}$ is strongly $r$-local for some $r$ depending only on $\tau$.
Without loss of generality, we assume that ${\cal C}$ is closed under the operation of adding a leaf to an arbitrary vertex. The class ${\cal C}$ still admits product structure, since we can simply add leaves to the factor $M$ of Definition 2.1.
Claim 13.
A class ${\cal C}$ closed under the operation of adding a leaf to an arbitrary vertex does not need coloring nor copying.
We include a folklore proof for the sake of completeness.
Proof 3.18 (Subproof).
First, we show how to “simulate” $k$-copying by adding leaves and colors.
Let $G\in{\cal C}$ be arbitrary. We denote by $kG$ its $k$-copy on vertex set $V(kG)=[k]\times V(G)$. Let $H$ be a graph obtained from $G$ by adding $k$ leaves $w_{1}^{v},w_{2}^{v},\ldots,w_{k}^{v}$ to each vertex $v\in V(G)$. Consider a coloring $c$ in which all vertices of $V(G)\subseteq V(H)$ have color 00, and each vertex $w_{i}^{v}$ has color $i$. Using $c$-colored $H$, we interpret a graph $H^{+}$ such that $H^{+}\left[\bigcup_{v\in V(G)}\{w_{1}^{v},\ldots,w_{k}^{v}\}\right]$ is isomorphic to $kG$ as witnessed by an isomorphism mapping vertex $w_{i}^{v}\in V(H^{+})$ to vertex $[i,v]\in V(kG)$.
It suffices to observe that $[i,v]$ and $[j,u]$ are adjacent in $kG$ if and only if $i=j$ and $uv\in E(G)$, that is, if and only if $w_{i}^{v}$ and $w_{j}^{u}$ have the same color $c(w_{i}^{v})=c(w_{j}^{u})>0$ and there is a path $(w_{i}^{v},a,b,w_{j}^{u})$ of length $3$ between $w_{i}^{v}$ and $w_{j}^{u}$ (because the only neighbor $a$ of $w_{i}^{v}$ is $v$, resp. $b$ of $w_{j}^{u}$ is $u$). This is clearly FO expressible.
Two vertices $w_{i}^{v}$ and $w_{j}^{u}$ are the copy of the same vertex $v=u$ if and only if $w_{i}^{v}$ and $w_{j}^{u}$ share their unique neighbor, so we can FO-define the relation $SV$ (cf. the definition of the $k$-copy operation). Two vertices $w_{i}^{v}$ and $w_{j}^{u}$ belong to the same copy of $G$ if and only if they have the same color $c(w_{i}^{v})=c(w_{j}^{u})\geq 1$, so we can FO-define the relation $SC$ (cf. definition of $k$-copy operation).
Thus far, we have shown how to transduce (without copying) a $k$-copy of any graph $G\in{\cal C}$ in another graph $H\in{\cal C}$.
As the second and final step of this proof, we show how to “simulate” coloring.
Let $G\in{\cal C}$ be an arbitrary graph and let $c:V(G)\to[k]$ be its arbitrary $k$-coloring (where $k\geq 1$). Let $H$ be a graph obtained from $G$ by adding $c(v)+1$ leaves $w_{0}^{v},w_{1}^{v},\ldots,w_{k}^{v}$ to each vertex $v\in V(G)$. Observe that, $v\in V(H)$ is a vertex of $V(G)\subseteq V(H)$ if and only if the degree of $v$ in $H$ is at least 2. Furthermore, $v\in V(G)\subseteq V(G)$ has color $c(v)$ if and only if $v$ has exactly $c(v)+1$ neighbors of degree one in $H$. Both of these properties are FO-expressible, so we can interpret the $c$-colored graph $G$ is the graph $H\in{\cal C}$.
Altogether, we get that for any transduction $\tau$ there is an interpretation $\iota$ such that $\tau({\cal C})\subseteq\iota({\cal C})$.
Using 13, we observe that we can modify $\xi^{\prime}(x_{1},x_{2})$ to obtain a formula $\xi(x_{1},x_{2})$ such that, for every $G^{\prime}\in C({\cal C})$ and every coloring $c$ of $G^{\prime}$ using the colors referenced in $\xi^{\prime}$, there is a graph $G^{\prime\prime}\in{\cal C}$ such that, for every $u,v\in V(G^{\prime})$, we have $G^{\prime\prime}\models\xi^{\prime}(u,v)$ iff $G^{\prime},c\models\xi(u,v)$.
Hence, it suffices to show, for every strongly local first-order formula $\xi$, that the $\xi$-interpretation of a graph $G\in{\cal C}$ has bounded ${\cal Q}^{\circ}_{r}$-clique-width. Since $G\subseteq Q\boxtimes M$ (Definition 2.1) for some $Q\in{\cal Q}$ and some graph $M$ of bounded tree-width, we apply Lemma 3.1 onto the following objects:
  • the 2-coloring $H$ of the graph $Q\boxtimes M$, where the color of $v\in V(H)$ is 1 if $v\in V(G)$ and 2 otherwise,
  • and the formula $\xi_{1}$ created from $\xi$ by restricting each quantifier to vertices of color 1 (i.e., belonging to $G$).
Doing so, we obtain a $(Q^{\circ}_{r},\ell)$-expression $\phi$ for a graph containing $\xi(G)$ as an induced subgraph, where $\ell$ depends on $\tau$, the degree bound of ${\cal Q}$ and the tree-width of $M$, and $Q^{\circ}_{r}\in{\cal Q}^{\circ}_{r}$.
Thus, we have shown that $\tau({\cal C})$ is a perturbation of the class ${\cal D}=\{\xi(G)|\>G\in{\cal C}\}$ of bounded ${\cal Q}^{\circ}_{r}$-clique-width.
In the special case of paths, we can strengthen Theorem 3.16. Let ${\cal P}^{\circ}$ denote the class of reflexive closures of all paths.
Theorem 3.19.
Let ${\cal C}$ be a class of graphs admitting product structure. Let $\tau$ be a first-order transduction such that $\tau({\cal C})$ is a class of simple graphs. Then, the class $\tau({\cal C})$ is a perturbation of a class of bounded ${\cal P}^{\circ}$-clique-width.
Proof 3.20.
From Theorem 3.16 we get that $\tau({\cal C})$ is a perturbation of a class of bounded ${\cal P}_{r}^{\circ}$-clique-width for some constant $r$, where ${\cal P}_{r}^{\circ}$ is the class of reflexive closures of $r$-th powers of paths.
We show that, given any $(P_{r}^{\circ},\ell)$-expression $\psi$ valued graph $G$, where $P_{r}^{\circ}\in{\cal P}_{r}^{\circ}$, one can build a $(Q^{\circ},3r\cdot\ell)$-expression $\phi$ valued $G$, where $Q^{\circ}\in{\cal P}^{\circ}$.
Let $(p_{0},p_{1},\ldots,p_{n})$ be the vertices of $P_{r}^{\circ}$ in the natural order along the underlying path of $P_{r}^{\circ}$. Observe that contracting each vertex $p_{i}$ to $p_{\left\lfloor\frac{i}{r}\right\rfloor r}$ creates a reflexive path $Q^{\circ}$ from $P_{r}^{\circ}$. We obtain a $(Q^{\circ},3r\cdot\ell)$-expression $\phi$ valued $G$ from $\psi$ by replacing each parameter vertex $p_{i}$ by $p_{\left\lfloor\frac{i}{r}\right\rfloor r}$, and using $3r$-times more colors to emulate the behavior of edge addition in $\phi$.
More specifically, in each subexpression $\alpha$ of $\phi$ corresponding to a subexpression $\beta$ of $\psi$, the color of a vertex $v$ is $(c_{v},i_{v}\mod 3r)$, where $i_{v}$ is the index of the parameter vertex $p_{i_{v}}$ of $v$ and $c_{v}$ is the color of $v$ in $\beta$. Then, each recoloring operation in $\psi$ corresponds to a sequence of recoloring operations in $\phi$, disjoint union operations remain unchanged, each operation creating a new vertex uses the modified color, and each edge addition operation between colors $c^{1}$ and $c^{2}$ is emulated as follows. For each pair $i,j$ such that $|i-j|\equiv 1\mod 3r$ or $|i-j|\equiv 0\mod 3r$, we create edges between vertices of colors $(c^{1},i)$ and $(c^{2},j)$ provided their parameter vertices are adjacent. We observe that, by doing so, we create in $\phi$ exactly the same edges as in $\psi$.
Remark 3.21.
The proof of Lemma 3.1 is constructive. We can interpret this proof as an FPT-time algorithm for computing a $(Q_{r}^{\circ},\ell)$-expression for graphs created by a $\xi$-interpretation of a 2-coloring of $Q\boxtimes M$, parameterized by $\xi$, the tree-width of $M$, and the maximum degree of $Q$; provided that $\xi$, $Q$, $M$, and the 2-coloring of $Q\boxtimes M$ are a part of the input.
Note, on the other hand, that knowing the 2-coloring of $Q\boxtimes M$ is crucial for an efficient computation of our $(Q_{r}^{\circ},\ell)$-expression since, in particular, already testing embeddability of a graph in the strong product of a path and a graph of bounded tree-width (tree-depth) is NP-hard [DBLP:conf/cccg/BiedlEU23].
Consequently, for Theorems 3.16 and 3.19, we also get analogous FPT-time algorithms for computing $(Q_{r}^{\circ},\ell)$- or $(P^{\circ},\ell)$-expressions of the members of $\tau({\cal C})$, but, importantly, for computing an expression for a graph $G^{\prime}\in\tau(G)$ where $G\in{\cal C}$, the algorithms based on Lemma 3.1 require $G$ and the particular coloring used by the transduction $\tau$ on the input.
On the other hand, by analyzing the proof of Theorem 2.5 in [DBLP:conf/csl/NesetrilMS22], we observe that the “splitting” of a transduction $\tau$ into a composition of a perturbation, an immersive transduction, and copying can be computed in FPT time on classes admitting FPT FO model checking, such as on considered ${\cal C}$, provided that the input contains $\tau$ and $G\in{\cal C}$ together with its coloring.
4 Backwards translation
In regard of Theorem 3.16 it is natural to ask to which extent the converse direction may hold; that is, whether each class being a perturbation of a class of bounded ${\cal Q}^{\circ}$-clique-width is a transduction of a class admitting a ${\cal Q}$-product structure. This, however, cannot be true in full generality due to the following case (with ${\cal Q}^{\circ}$ containing only the single-vertex loop):
Theorem 4.1 (Nešetřil, Ossona de Mendez, Pilipczuk, Rabinovich, and Siebertz [DBLP:conf/soda/NesetrilMPRS21]).
If a class of graphs ${\cal C}$ has bounded clique-width,444[DBLP:conf/soda/NesetrilMPRS21] formulated the statement with functionally equivalent rank-width. then the following conditions are equivalent:
  • ${\cal C}$ has a stable edge relation,
  • ${\cal C}$ is a transduction of a class of bounded tree-width.
On the other hand, adding a suitable stability assumption already allows to formulate the converse in full strength.
Theorem 4.2.
Let ${\cal Q}$ be a class of bounded-degree graphs, let ${\cal Q}^{\circ}$ be the reflexive closure of ${\cal Q}$, and let $k$ be an integer. Let ${\cal C}$ be a graph class such that, for each $G\in{\cal C}$, there is a $k$-stable $(Q^{\circ},k)$-expression $\phi$ for some $Q^{\circ}\in{\cal Q}^{\circ}$, such that $G$ is a $k$-perturbation of the value of $\phi$.
Then, there is a class ${\cal D}$ admitting ${\cal Q}$-product structure and a first-order transduction $\tau$ such that ${\cal C}\subseteq\tau({\cal D})$.
Before we prove Theorem 4.2, we remark that its view of stability is equivalent to the more traditional view as follows:
Claim 14.
Let ${\cal C}$ be a class of graphs and let ${\cal Q}$ be a class of bounded degree graphs and let ${\cal Q}^{\circ}$ be the reflexive closure of ${\cal Q}$. Then, the following are equivalent:
  • (A)
    There are constants $k$ and $\ell$ such that, for each graph $G\in{\cal C}$, there is a graph $Q^{\circ}\in{\cal Q}^{\circ}$ and a $k$-stable $(Q^{\circ},\ell)$-expression valued $G$.
  • (B)
    There is a stable bounded clique-width class ${\cal D}$ such that, for each graph $G\in{\cal C}$, there are graphs $H\in{\cal D}$ and $Q\in{\cal Q}$ such that $G$ is an induced subgraph of $Q\boxtimes H$.
Proof 4.3 (Subproof).
We show that (A) implies (B). Let $Q\in{\cal Q}$ be a graph and let $Q^{\circ}\in{\cal Q}$ be its reflexive closure. Let $\phi$ be a $k$-stable $(Q^{\circ},\ell)$-expression valued $G$, and let $\phi_{depar}$ be its deparameterization valued $G_{depar}$. Then, $G_{depar}$ has clique-width at most $\ell$, and $G_{depar}$ excludes bi-induced half-graph of order $k$. Using the construction from the proof of Theorem 2.4, we observe that $G$ is an induced subgraph of $Q\boxtimes G_{depar}$.
Let ${\cal D}:=\{G_{depar}|\>G\in{\cal C}\}$ be the class obtained from ${\cal C}$ by the above described process. Then, it follows from Theorem 4.1 that ${\cal D}$ is stable. Furthermore, its clique-width is at most $\ell$, and so ${\cal D}$ witnesses that (B) holds.
We show that (B) implies (A). Let $G\in{\cal C}$ be arbitrary. Let $H\in{\cal D}$ and $Q\in{\cal Q}$ be graphs satisfying $G\subseteq_{i}Q\boxtimes H$. Let $Q^{\circ}$ be the reflexive closure of $Q$. Let $\ell$ be the clique-width of ${\cal D}$, and let $k^{\prime}$ be the order of the largest bi-induced half-graph in ${\cal D}$.
Let $\phi$ be the $(Q^{\circ},\ell)$-expression valued $G$ constructed in the proof of Theorem 2.4. We observe that, for each $h\in V(H)$, there is a subexpression $\phi_{h}$ valued the copy $G[V(Q)\boxtimes\{h\}]$ of $Q$. Furthermore, all vertices in $\phi_{h}$ have the same color and form a clique in the value $G_{depar}$ of the deparameterization $\phi_{depar}$ of $\phi$, so they are twins in $G_{depar}$. That is, one can obtain $H$ from $G_{depar}$ by identifying twins, and so any induced subgraph without twins in $G_{depar}$ is also an induced subgraph of $H$.
Suppose that there is a bi-induced half-graph $J$ in $G_{depar}$ of order $k^{+}:=2(k^{\prime}+2)^{2}$. We obtain a contradiction by finding a bi-induced half-graph of order $k^{\prime}+1$ in $H$:
Let $A=(a_{1},a_{2},\ldots,a_{k^{+}}),B=(b_{1},b_{2},\ldots,b_{k^{+}})$ be the natural bipartition of $J$ with the orderings on $A$ and $B$. That is, $a_{i}b_{j}$ is an edge if and only if $i\leq j$. Observe that, within each part $A$ or $B$, there are no twins, and so the possible twin pairs within $J$ necessarily form a matching between $A$ and $B$ (not perfect in general). Our aim thus is to find a bi-induced half-graph $J^{\prime}\subseteq J$ of order $k^{\prime}+1$ which avoids all these twin pairs.
For $i\in[2k^{\prime}+4]$, we denote by $A_{i}:=\{a_{j}|$ $(i-1)(k^{\prime}+2)+1\leq j\leq i(k^{\prime}+2)\}\subseteq A$ and $B_{i}:=\{b_{j}|$ $(i-1)(k^{\prime}+2)+1\leq j\leq i(k^{\prime}+2)\}\subseteq B$. Next, for $j=1,2,\ldots,k^{\prime}+1$ in this order, we choose (arbitrarily) $a_{i_{j}}\in A_{2j-1}$ such that $a_{i_{j}}$ has no twin among the previously chosen vertices of $B$ – this is possible since $|A_{2j-1}|=k^{\prime}+2>j$, and we analogously choose $b_{i^{\prime}_{j}}\in B_{2j+2}$ such that $b_{i^{\prime}_{j}}$ has no twin among the previously chosen vertices of $A$. Then $J^{\prime}:=J[\bigcup_{j\in[k^{\prime}+1]}\{a_{i_{j}},b_{i^{\prime}_{j}}\}]$ is a bi-induced half-graph of order $k^{\prime}+1$ in $G_{depar}$. Furthermore, $J^{\prime}$ has no twins, so it is isomorphic to an induced subgraph of $H\in{\cal D}$, a contradiction.
Proof 4.4 (Proof of Theorem 4.2).
Let $G\in{\cal C}$ be arbitrary. Let $Q_{G}^{\circ}\in{\cal Q}^{\circ}$, and let $\phi_{G}$ be a $k$-stable $(Q_{G}^{\circ},k)$-expression such that $G$ is a $k$-perturbation of the value $G^{\prime}$ of $\phi_{G}$. We denote by $Q_{G}^{\prime}\in{\cal Q}$ the graph obtained from $Q_{G}^{\circ}$ by removing all loops. Let $\psi_{G}$ be the deparameterization of $\phi_{G}$. We denote by $M_{G}$ the value of $\psi_{G}$, which is of clique-width at most $k$. Since $\phi_{G}$ is assumed $k$-stable, $M_{G}$ does not contain bi-induced half-graph of order $k$.
Consider the class =${\cal M}=\{M_{G}|\>G\in{\cal C}\}$ which, by the previous, has stable edge relation. So, by Theorem 4.1, there is a first-order transductions $\hat{\tau}$ and a class $\hat{{\cal M}}$ of graphs of bounded tree-width such that $^{\cal M}\subseteq\hat{\tau}(\hat{{\cal M}})$.
Let ${\cal M}^{tw}$ be the closure of $\hat{{\cal M}}$ under the operation of adding a leaf to an arbitrary vertex. The tree-width of ${\cal M}^{tw}$ is still bounded, and the class ${\cal M}^{tw}$ does not need copying nor coloring. So, there is an interpretation $\tau^{\prime}$ such that ${\cal M}\subseteq\tau^{\prime}({\cal M}^{tw})$. Let $\xi$ be the binary formula used in $\tau^{\prime}$ for the edge relation.
Notably, there is a graph $M_{G}^{tw}\in{\cal M}^{tw}$ such that $M_{G}$ is an induced subgraph of $\tau^{\prime}(M_{G}^{tw})$. In order to simplify our notation, we assume that $V(M_{G})\subseteq V(M_{G}^{tw})$ with identity map of vertices, and so that for any pair of vertices $u,v\in V(M_{G})$ we have $uv\in E(M_{G})$ if and only if $M_{G}^{tw}\models\xi(u,v)$.
We denote by $M_{G}^{uni}$ the graph obtained from $M_{G}^{tw}$ by adding a universal vertex $u_{G}^{uni}$, which increases the tree-width by at most one. Our core step is to transduce, with a suitable coloring, $Q^{\prime}_{G}\boxtimes M_{G}$ from $Q^{\prime}_{G}\boxtimes M_{G}^{uni}$ using a transduction formula independent of a particular choice of $G$.
Claim 15.
Let $X=V\big{(}Q^{\prime}_{G}\boxtimes M_{G}^{uni}\big{)}$. There exist first-order formulas (constructed independently of $G$) which, for a suitable coloring of $X$, express the following properties of any vertex pair $x,y\in X$ over $Q^{\prime}_{G}\boxtimes M_{G}^{uni}$:
  1. (i)
    that $x=[q,m]$ and $y=[q^{\prime},m^{\prime}]$ for $q=q^{\prime}$ (“same row”),
  2. (ii)
    that $x=[q,m]$ and $y=[q^{\prime},m^{\prime}]$ for $qq^{\prime}\in E(Q^{\prime}_{G})$,
  3. (iii)
    assuming that $x=[q,m]$ and $y=[q^{\prime},m^{\prime}]$ satisfy $q=q^{\prime}\vee qq^{\prime}\in E(Q^{\prime}_{G})$, the properties that $m=m^{\prime}$ (“same column”) and that $mm^{\prime}\in E(M_{G}^{tw})$.
Proof 4.5 (Subproof).
A coloring of a graph $G$ is proper if every edge of $G$ receives distinct colors for its ends. Since the factor $Q^{\prime}_{G}$ is of bounded degree and the factor $M_{G}^{uni}$ is degenerate thanks to being of bounded tree-width, there exist proper colorings $c_{1}$ and $c_{2}$ of $Q^{\prime}_{G}$ and $M_{G}^{uni}$, respectively, each using a bounded number of colors. We color $X$ by the Cartesian product $c=c_{1}\times c_{2}$ of the two colorings. We proceed to define the formulas expressing (i), (ii), and (iii):
(i) The formula $\sigma_{1}$ reads that $c_{1}(x)=c_{1}(y)$ and $xy$ is an edge, or that there exists a vertex $z$ such that $c_{1}(x)=c_{1}(y)=c_{1}(z)$ and $xz,zy$ are edges.
Assume $Q^{\prime}_{G}\boxtimes M_{G}^{uni}\models\sigma_{1}(x,y)$. If $xy$ is an edge, but $q\not=q^{\prime}$, then $qq^{\prime}\in E(Q^{\prime}_{G})$ by the definition of a strong product. However, $c_{1}(x)=c_{1}(y)$ is in a contradiction with $c_{1}$ being a proper coloring of $Q^{\prime}_{G}$. On the other hand, if $xz,zy$ are edges where $z=[q^{\prime\prime},m^{\prime\prime}]$, and $q\not=q^{\prime}$, then, up to symmetry, we get $q\not=q^{\prime\prime}$. As in the previous subcase, we have $c_{1}(x)=c_{1}(z)$ and $qq^{\prime\prime}\in E(Q^{\prime}_{G})$, which is again a contradiction to $c_{1}$.
Now assume $q=q^{\prime}$. Therefore, trivially $c_{1}(x)=c_{1}(y)$. We choose $z$ as $z=[q,u_{G}^{uni}]$ where $u_{G}^{uni}$ is the universal vertex. If $z=x$ (up to symmetry), we have $m^{\prime}u_{G}^{uni}\in E(M_{G}^{uni})$, and so $zy=xy$ is an edge by the definition of a strong product. If $z\not\in\{x,y\}$, then we likewise get that both $xz$ and $zy$ are edges. In both cases, $Q^{\prime}_{G}\boxtimes M_{G}^{uni}\models\sigma_{1}(x,y)$.
(ii) Using the formula of (i), this formula $\sigma_{2}$ reads that $x$ and $y$ are not from the same row, and there exists $z$ such that $z$ is in the same row with $x$ and $yz$ is an edge.
Assume $Q^{\prime}_{G}\boxtimes M_{G}^{uni}\models\sigma_{2}(x,y)$. Then $q\not=q^{\prime}$ and $z=[q,m^{\prime\prime}]$, and $qq^{\prime}\in E(Q^{\prime}_{G})$ by the definition of a strong product since $yz$ is an edge. Conversely, assume $qq^{\prime}\in E(Q^{\prime}_{G})$, which also means $q\not=q^{\prime}$. We choose $z=[q,m^{\prime}]$ (same column as $y$ and same row as $x$), and hence $yz$ is an edge by the definition of a strong product. Hence, $Q^{\prime}_{G}\boxtimes M_{G}^{uni}\models\sigma_{2}(x,y)$.
(iii) We first, under the assumption that $x=[q,m]$ and $y=[q^{\prime},m^{\prime}]$ satisfy $q=q^{\prime}\vee qq^{\prime}\in E(Q^{\prime}_{G})$, express that $m=m^{\prime}$. We give a formula $\sigma_{3}$ reading that $x=y$, or $xy$ is an edge and $c_{2}(x)=c_{2}(y)$. We claim that, under the stated assumptions on $x$ and $y$$Q^{\prime}_{G}\boxtimes M_{G}^{uni}\models\sigma_{3}(x,y)$ if and only if $m=m^{\prime}$. Indeed, as in part (i), the third possibility of $mm^{\prime}\in E(M_{G}^{tw})$ is now excluded by $c_{2}$ being a proper coloring of $M_{G}^{uni}$.
The last formula $\sigma_{4}$ then reads that $x$ and $y$ are not in the same column, $c_{2}(x),c_{2}(y)\not=c_{2}(u_{G}^{uni})$, and there exists $z$ such that $z$ is in the same column with $x$ and $yz$ is an edge. The conclusion that $Q^{\prime}_{G}\boxtimes M_{G}^{uni}\models\sigma_{4}(x,y)$ $\iff$ $mm^{\prime}\in E(M_{G}^{tw})$ under the stated assumptions then follows in the same way as in part (ii).
Observe that the formulas of (iii), with a bit of technical arguments, allows us to “see” the factor $M_{G}^{tw}$ within the product $Q^{\prime}_{G}\boxtimes M_{G}^{uni}$ and, consequently, to express the result of the interpretation $\tau^{\prime}(M_{G}^{tw})$. The latter, in turn, together with (i) and (ii) define the edge set of $Q^{\prime}_{G}\boxtimes M_{G}$, as desired. Finally, we use Theorem 2.4 to argue that an induced subgraph of $Q^{\prime}_{G}\boxtimes M_{G}$ is the value of the $(Q_{G}^{\circ},k)$-expression $\phi_{G}$ we started with.
Formally, our proof is finished as follows. In view of our assumptions, namely $V(M_{G})\subseteq V(M_{G}^{uni})$, it suffices to construct a binary first-order formula $\varrho$ such that $Q^{\prime}_{G}\boxtimes M_{G}^{uni}\models\varrho(x,y)$ $\iff$ $xy\in E\big{(}Q^{\prime}_{G}\boxtimes M_{G}\big{)}$. Let $x=[q,m]$ and $y=[q^{\prime},m^{\prime}]$.
If $q\not=q^{\prime}$ and $qq^{\prime}\not\in E(Q^{\prime}_{G})$, which is expressed using 15 – formulas (i) and (ii), then $\varrho$ states that $xy$ is not an edge. This is correct by the definition of a strong product. Hence, for the remaining parts of the formula $\varrho$, we may assume that $q=q^{\prime}$ or $qq^{\prime}\in E(Q^{\prime}_{G})$ as in 15 (iii).
Under the latter assumption, we first show how to interpret from the graph $Q^{\prime}_{G}\boxtimes M_{G}^{uni}$ a graph $M_{x,y}$ isomorphic to $M_{G}^{tw}$ such that, moreover, $x,y\in V(M_{x,y})$. This is done as follows. Let $Y_{0}=\{q\}\times V(M_{G}^{tw})$ and $Y=(Y_{0}\setminus\{[q,m^{\prime}]\})\cup\{y\}$. We choose $Y$ as the domain (vertex set) of $M_{x,y}$, which can be expressed using the formulas constructed in 15 and the unique color $c_{2}(u_{G}^{uni})$ of the universal vertex $u_{G}^{uni}$ of $M_{G}^{uni}$. We likewise interpret the edge relation of $M_{G}^{tw}$ using specifically 15 (iii), which finishes our interpretation $\iota$.
Now, we recall the interpretation $\tau^{\prime}$ such that $M_{G}$ is an induced subgraph of $\tau^{\prime}(M_{G}^{tw})$, and the fact that $\tau^{\prime}$ does not use coloring. Hence, we can compose $\tau^{\prime}$ after $\iota$ (since $\tau^{\prime}(M_{G}^{tw})$ is isomorphic to $\tau^{\prime}(M_{x,y})$) to interpret the edge relation of $M_{G}$. The latter, together with the formulas of 15, define the sought formula $\varrho$.
Finally, since the class $\{Q^{\prime}_{G}\boxtimes M_{G}^{uni}|\>G\in{\cal C}\}$ admits ${\cal Q}$-product structure by the definition, and the hereditary closure can be expressed as a transduction, too, the only remaining task is to argue that the value of the $(Q_{G}^{\circ},k)$-expression $\phi_{G}$ we started with is indeed an induced subgraph of $Q^{\prime}_{G}\boxtimes M_{G}$. Examining the (short) proof of Theorem 2.4 in [DBLP:conf/mfcs/HlinenyJ24], one may find out that the bounded-clique-width factor used there is actually obtained as the value of the deparameterization of $\phi_{G}$, as we do here with $M_{G}$. The whole proof is finished.
5 Excluding grids in transductions
Our last contribution are two example application of the results of Section 3; showing that
  • the class of all 3D grids, and
  • the class of certain non-local modifications of 2D grids,
are not transducible from any class admitting product structure, and in particular not from the class of planar graphs.
Recall that ${\cal P}^{\circ}$ denotes the class of reflexive closures of paths.
Theorem 5.1.
For any $k$ and $n$ positive integers, the ${\cal P}^{\circ}$-clique-width of any $k$-perturbation of an $n\times n\times n\,$-grid is in $\Omega_{k}(n)$.
Proof 5.2.
Let $G$ be our $n\times n\times n$-grid. Let $H$ be any $k$-perturbation of $G$. Let ${\cal L}=(L_{1},L_{2},\ldots,L_{k})$ be a partition of $V(H)=V(G)$ and $\sim$ a symmetric relation on $[k]$ witnessing that $H$ is a $k$-perturbation of $G$. We say that a pair of parts $(L_{i},L_{j})$ is flipped if $i\sim j$, otherwise we say that it is not flipped.
Consider a subgraph $G^{\prime}$ of $G$. A part $L_{i}\in{\cal L}$ is $V(G^{\prime})$-empty if $V(G^{\prime})\cap L_{i}=\emptyset$ and $V(G^{\prime})$-small if $|V(G^{\prime})\cap L_{i}|\leq 12$. We use the following simple claim:
Claim 16.
Let $G_{1}\subseteq_{i}G$ be a grid of size $m^{\prime}\times m^{\prime}\times m^{\prime}$, where $m^{\prime}\geq 3m^{\prime\prime}$. If a part $L_{i}\in{\cal L}$ is $V(G_{1})$-small, then there exists an induced subgrid $G_{2}\subseteq_{i}G_{1}$ of size $m^{\prime\prime}\times m^{\prime\prime}\times m^{\prime\prime}$ such that $L_{i}$ is $V(G_{2})$-empty.
Proof 5.3 (Subproof).
At most 12 of $273^{3}=$ 27 pairwise disjoint subgrids of $G_{1}$ of size $m^{\prime\prime}\times m^{\prime\prime}\times m^{\prime\prime}$ may intersect $L_{i}$.
We inductively construct a sequence of graphs $G^{\prime}_{0}:=G,$ $G^{\prime}_{1},\ldots,G^{\prime}_{\ell}$ such that, for each $i\geq 0$, if some part of ${\cal L}$ is $V(G^{\prime}_{i})$-small but not $V(G^{\prime}_{i})$-empty, then let $G^{\prime}_{i+1}$ be equal to the graph $G_{2}$ of 16 called with $G_{1}=G^{\prime}_{i}$. Then every part of ${\cal L}$ is either $V(G^{\prime}_{\ell})$-empty, or not $V(G^{\prime}_{\ell})$-small.
Moreover, since every step $i$ in the construction of $G^{\prime}_{\ell}$ adds a new $V(G^{\prime}_{i+1})$-empty part, we have $\ell\leq k$, and so $G^{\prime}_{\ell}$ is an induced subgrid of $G$ of size $m\times m\times m$ where $m\geq\lfloor n/3^{k}\rfloor$, that is, $m=\Theta_{k}(n)$. We denote by $G^{\prime}:=G^{\prime}_{\ell}$ and by ${\cal L}^{\prime}$ the restriction of ${\cal L}$ to $V(G^{\prime})$.
We observe that, there is a symmetric relation $\sim^{\prime}$, such that ${\cal L}^{\prime}$ and $\sim^{\prime}$ witness that $H^{\prime}:=H[V(G^{\prime})]$ is a $k^{\prime}$-perturbation of $G^{\prime}$ for some $k^{\prime}\leq k$. Note that each nonempty part of ${\cal L}^{\prime}$ has size at least $13$.
Observe that the $m\times m\times m$ grid $G^{\prime}$ has a maximum degree at most 6 and the diameter less than $3m$. We show that $H^{\prime}$ has diameter less than $9m$, which follows from:
Claim 17.
If $u$ and $v$ are neighbors in $G^{\prime}$, then the $u$$v$ distance in $H^{\prime}$ is at most $3$.
Proof 5.4 (Subproof).
Let $L_{i}$ be the part containing $u$ and let $L_{j}$ be the part containing $v$. If $(L_{i},L_{j})$ is not flipped, then the claim is trivial. Hence, we assume that $(L_{i},L_{j})$ is flipped. Since $u$ has at most $6$ $G^{\prime}$-neighbors in $L_{j}$, all but at most $6$ vertices of $L_{j}$ are neighbors of $u$ in $H^{\prime}$, that is, $u$ has at least $7$ $H^{\prime}$-neighbors in $L_{j}$. We analogously obtain that $v$ has an $H^{\prime}$-neighbor $w$ in $L_{i}$, and all but at most $6$ vertices of $L_{j}$ are neighbors of $w$ in $H^{\prime}$. Hence, by pigeon-hole principle, $u$ and $w$ have common $H^{\prime}$-neighbor in $L_{j}$, or $w$ is an $H^{\prime}$-neighbor of $u$ if $i=j$, and so the distance between $u$ and $v$ in $H^{\prime}$ is at most $3$.
Let $\ell$ be the smallest number such that there is a reflexive path $P\in{\cal P}^{\circ}$ and a $(P,\ell)$-expression $\phi$ valued $H$. Since $H^{\prime}$ is an induced subgraph of $H$, there is a $(P^{\prime},\ell)$-expression $\phi^{\prime}$ valued $H^{\prime}$ for some reflexive subpath $P^{\prime}\in{\cal P}^{\circ}$ of $P$.
We assume that $P^{\prime}$ is the shortest possible such path. Then, the natural homomorphism from $H^{\prime}$ to $P^{\prime}$ is surjective, so $P^{\prime}$ has at most $9m$ vertices because the diameter of $H^{\prime}$ is at most $19m-$ 1. Let $p_{1},p_{2},\ldots,p_{9m}$ be the vertices of $P^{\prime}$ in the natural order along the path.
The following claim is folklore, and we include a proof for the sake of completeness:
Claim 18.
There is a subexpression $\alpha$ of $\phi^{\prime}$ such that $\alpha$ is the disjoint union of two smaller subexpressions $\alpha_{1}$ and $\alpha_{2}$, such that $\alpha_{1}$ contains between $\frac{1}{3}m^{3}$ and $\frac{2}{3}m^{3}$ vertices.
Proof 5.5 (Subproof).
We can see $\phi^{\prime}$ as a rooted binary tree $T^{\prime}$ in which the nodes with two children correspond to the disjoint union operations, leafs correspond to the operations creating a new vertex, and nodes with one child correspond to remaining operations. We find the node corresponding to a sought subexpression $\alpha$ as the bottommost union node $a$ of $T^{\prime}$ such that $\alpha$ contains at least $\frac{2}{3}m^{3}$ vertices. Then, since $a$ has two children, in one of the two subexpressions there has to be at least $\frac{1}{2}\cdot\frac{2}{3}m^{3}=\frac{1}{3}m^{3}$ vertices (and less than $\frac{2}{3}m^{3}$).
Hence, there is a partition $(A,B)$ of $V(H^{\prime})=V(G^{\prime})$, such that vertices belonging to $A$ are exactly those contained within $\alpha_{1}$, and $B$ contains the remaining vertices. Notably, the size of both $A$ and $B$ is between $\frac{1}{3}m^{3}$ and $\frac{2}{3}m^{3}$. From this property, we can already derive the following important claim:
Claim 19.
In the $m\times m\times m\,$-grid $G^{\prime}$, there is an induced matching $M$ between $A$ and $B$ of size at least $\frac{1}{36\cdot 61}m^{2}$.
Proof 5.6 (Subproof).
We shortly refer to the parts $A$ and $B$ as to colors of $V(G^{\prime})$; that is, $v\in V(G^{\prime})$ is of color $A$ if $v\in A$, and color $B$ otherwise. In the proof we are going to find at least $\frac{1}{36}m^{2}$ 2-colored edges in $G^{\prime}$. Then, since the maximum degree in $G^{\prime}$ is $6$, we can always greedily pick a $\frac{1}{61}$-fraction of them to form the sought induced matching $M$ – each picked edge $uw$ “excludes” up to $5$ edges incident to $u$, up to $5$ edges incident to $w$, and up to $5\cdot(5+5)=50$ more edges incident to the previous edges.
The grid $G^{\prime}$ can be clearly vertex-partitioned into $m$ copies of the $m\times m\,$-grid, which we further call the planes of $G^{\prime}$, and each plane $K$ can be similarly partitioned into copies of the $m$-vertex path as into the rows of $K$ or into the columns of $K$, further called lines of $G^{\prime}$. A set $X\subseteq V(G^{\prime})$ is monochromatic if $X\subseteq A$ or $X\subseteq B$, and $X$ is $\varepsilon$-balanced if at least $\varepsilon\cdot|X|$ vertices of it belong to each color $A$ and $B$. Observe that every non-monochromatic line of $G^{\prime}$ contains a 2-colored edge.
In the first step we show that every $\varepsilon$-balanced, $\varepsilon\leq\frac{1}{3}$, plane $K$ of $G^{\prime}$ contains $\geq\varepsilon m\>$ 2-colored edges. If there are disjoint monochromatic lines in $K$ of opposite colors, then all perpendicular lines are non-monochromatic, and hence we already have $m$ 2-colored edges. Hence, all monochromatic lines, say rows, are of the same color ($A$ or $B$), and due to the $\varepsilon$-balanced condition, there are less than $(1-\varepsilon)m$ of them. So, we get $\geq\varepsilon m\>$ 2-colored edges from the remaining non-monochromatic rows.
In the second step, we finish by examining $\frac{1}{6}$-balanced planes of $G^{\prime}$. If there are planes $K_{1}$ and $K_{2}$ which are not $\frac{1}{6}$-balanced, and $K_{1}$ has majority of $A$-color while $K_{2}$ majority of $B$-color, then among the lines perpendicular to them at least $1-\frac{2}{6}=\frac{2}{3}$ fraction of them are non-monochromatic, giving us the needed number of 2-colored edges. Otherwise, all planes of $G^{\prime}$ which are not $\frac{1}{6}$-balanced, have majority in the same color, and so, by the property of $V(G^{\prime})$ being $\frac{1}{3}$-balanced, there are less than $\frac{5}{6}m$ of them. In the remaining at least $\frac{1}{6}m$ planes which are $\frac{1}{6}$-balanced, we find at least $\frac{1}{36}m^{2}$ 2-colored edges by the first step argument.
Let $M$ be the matching of size $|M|\geq\frac{m^{2}}{36\cdot 61}$ from 19. Recall that there are only $9m$ distinct parameter vertices from $P^{\prime}$. By the pigeon-hole principle, there is a parameter vertex $p_{i}$ such that at least $\frac{m^{2}}{36\cdot 61}/{9m}=\frac{m}{9\cdot 36\cdot 61}$ endpoints of $M$ belonging to $A$ have $p_{i}$ as their parameter vertex. We denote by $A^{M}_{i}$ these vertices, and we denote by $B^{M}_{i\pm 1}$ the vertices of $B$ adjacent to $A^{M}_{i}$ in $M$.
Observe that the only possible parameter vertices for $B^{M}_{i\pm 1}$ are $p_{i-1}$, $p_{i}$, and $p_{i+1}$. Take the expression $\psi$ obtained from $\phi^{\prime}$ by deleting all vertices $v$ whose parameter vertex $p$ is not one of $p_{i-1}$, $p_{i}$, and $p_{i+1}$. Let $\alpha^{\prime}$, $\alpha^{\prime}_{1}$, and $\alpha^{\prime}_{2}$ be the subexpressions of $\psi$ corresponding to $\alpha$, $\alpha_{1}$, and $\alpha_{2}$ of 18.
Suppose that there are at least $k^{2}+1$ vertices $u_{0},u_{1},\ldots,u_{k^{2}}\in A^{M}_{i}$ of the same color in (the value of) $\alpha_{1}^{\prime}$. Denote by $w_{0},w_{1},\ldots,w_{k^{2}}\in B^{M}_{i\pm 1}$ in order their neighbors within the matching $M$. By Definition 2.3, vertices of the same color and same parameter vertex in $\alpha^{\prime}_{1}$ must have the same adjacency to the vertices outside $\alpha^{\prime}_{1}$; in particular, to the vertices of $B^{M}_{i\pm 1}$. Moreover, by pigeon-hole principle, at least two of the vertices $u_{0},\ldots,u_{k^{2}}$, say $u_{1}$ and $u_{2}$ up to symmetry, belong to the same part of ${\cal L}^{\prime}$, and their mates $w_{1}$ and $w_{2}$ also belong to the same (another) part of ${\cal L}^{\prime}$. Hence, the vertices $u_{1}$ and $u_{2}$ have the same adjacency to $w_{1}$ and $w_{2}$, which is impossible in a matching or its flip.
Hence, vertices of $A^{M}_{i}$ use at least $\frac{1}{9\cdot 36\cdot 61k^{2}}\cdot m\in\Omega_{k}(n)$ distinct colors.
Corollary 5.7.
Let ${\cal C}$ be a graph class admitting a classical product structure. (Such as, ${\cal C}$ could be the class of planar graphs, or the class of graphs embeddable in a fixed surface.) Then, ${\cal C}$ does not transduce the class of all 3D grids.
Proof 5.8.
Suppose otherwise. Then, it follows from Theorem 3.19 that there is a graph class ${\cal G}$ of bounded ${\cal P}^{\circ}$-clique-width such that a perturbation of ${\cal G}$ contains the class of all 3D grids.
Since perturbations are involutory, we get that ${\cal G}$ is a perturbation of the class of all 3D grids, contradicting Theorem 5.1.
[…]
Theorem 5.9.
For any $k$ and $n$ positive integers, the ${\cal P}^{\circ}$-clique-width of any $k$-perturbation of the graph which is composed of $n$ disjoint copies of the pinned grid of order $n$, is in $\Omega_{k}(n)$.
Proof 5.10.
We follow similar ideas as in the proof of Theorem 5.1. Let $G=D_{1}\cup\ldots\cup D_{n}$ be the disjoint union of copies $D_{1},\ldots,D_{n}$ of the pinned grid $D$ of order $n$, and let $H$ be any $k$-perturbation of $G$ witnessed by the partition ${\cal L}=(L_{1},\ldots,L_{k})$ of $V(H)=V(G)$ and relation $\sim$ on $[k]$.
Let the vertices of $G$ which come as copies of the vertex $w$ of $D$ be called apex vertices of $G$ and the remaining be the grid vertices. Consider $G^{\prime}\subseteq G$. A part $L_{i}$ of ${\cal L}$ is $V(G^{\prime})$-small if $L_{i}$ contains at most $8$ grid vertices of $G^{\prime}$. A copy $D_{j}\subseteq G^{\prime}$ of $D$ is $V(G^{\prime})$-rare if the apex vertex $w_{j}$ of $D_{j}$ belongs to a part $w_{j}\in L_{i^{\prime}}\in{\cal L}$ such that $L_{i^{\prime}}$ contains no other apex vertex of $G^{\prime}$. We form $G^{\prime\prime}\subseteq G$ by iteratively removing from $G$ the copies $D_{j}$ of $D$ such that, at the respective iteration $G^{\prime}\subseteq G$,  $D_{j}$ contains a grid vertex of a $V(G^{\prime})$-small part of ${\cal L}$ or $D_{j}$ is $V(G^{\prime})$-rare. Observe that in this iterative process, every part of ${\cal L}$ can become small at most once, and every removal of a rare copy is counted towards a different part of ${\cal L}$. So, the whole process stops after at most $8k+k$ removal rounds, removing at most $9k$ copies of $D$ from $G$.
Consequently, whenever $n>9k$, we have got nonempty $G^{\prime\prime}\subseteq G$ and without loss of generality a pinned grid $D_{1}\subseteq G^{\prime\prime}$ of order $n$ such that; no $V(G^{\prime\prime})$-small part of ${\cal L}$ contains a grid vertex of $D_{1}$, and $D_{1}$ is not $V(G^{\prime\prime})$-rare. Let $H^{\prime\prime}\subseteq_{i}H$ denote the corresponding perturbation of $G^{\prime\prime}$. Let $w_{1}$ be the apex vertex of $D_{1}$. We get:
Claim 20.
Any two vertices of $V(D_{1})\setminus\{w_{1}\}$ have distance in $H^{\prime\prime}$ at most $418n+$ 4.
Proof 5.11 (Subproof).
Let $H^{\#}$ be $H^{\prime\prime}$ restricted to the grid vertices of $G^{\prime\prime}$, and $D_{1}^{\prime}:=D_{1}-w_{1}$. Since the degrees in 2D grids are $\leq 4$ and no small part of ${\cal L}$ intersects $V(D_{1}^{\prime})$, we analogously as in 17 derive that, for any $u,v\in V(D_{1}^{\prime})$ such that $uv\in E(D_{1}^{\prime})$, the $u$$v$ distance in $H^{\#}$ is at most $3$.
Let $W_{1}\subseteq V(D_{1}^{\prime})$ be the set of the $G^{\prime\prime}$-neighbors of $w_{1}$. Observe that any vertex of $D_{1}^{\prime}$ has $D_{1}^{\prime}$-distance $\leq 2n$ from $W_{1}$, and so $H^{\#}$-distance $\leq 6n$ from $W_{1}$ by the previous claim. Similarly, if we split $W_{1}$ into two nonempty parts, then the $H^{\#}$-distance between the two parts is $\leq 6n$, too. Pick any $u,v\in V(D_{1}^{\prime})$. If none of the edges incident to $w_{1}$ in $G^{\prime\prime}$ is flipped in $H^{\prime\prime}$, then there is a connection from $u$ to $v$ (via $w_{1}$) in $H^{\prime\prime}$ of length at most $26n+2+6n=12n+$ 2.
Further, since $D_{1}$ is not $V(G^{\prime\prime})$-rare, there is an apex vertex $w_{2}\in V(G^{\prime\prime})$, $w_{2}\not=w_{1}$, such that $\{w_{1},w_{2}\}\subseteq L_{i}\in{\cal L}$. Note that $w_{2}$ is not adjacent to any vertex of $D_{1}$ in $G^{\prime\prime}$. If all edges incident to $w_{1}$ in $G^{\prime\prime}$ are flipped in $H^{\prime\prime}$, then $w_{2}$ is adjacent to all vertices of $W_{1}$ in $H^{\prime\prime}$, and we are done as in the previous paragraph. Otherwise, every vertex of $W_{1}$ is in $H^{\prime\prime}$ adjacent to $w_{1}$ or $w_{2}$ (depending on which incident edges are flipped in $H^{\prime\prime}$). Consequently, there is a connection in $H^{\prime\prime}$ from $u$ via $w_{1}$ and $w_{2}$ (up to symmetry) to $v$ of total length at most $46n+2+6n+2+6n=18n+$ 4.
Continuing with the proof, let $\ell$ be the smallest number such that there is a reflexive path $P\in{\cal P}^{\circ}$ and a $(P,\ell)$-expression $\phi$ valued $H$. Since $H^{\prime\prime}$ is an induced subgraph of $H$, there is a $(P,\ell)$-expression $\phi^{\prime\prime}$ valued $H^{\prime\prime}$. Consider the $(P,\ell)$-expression $\phi_{1}$ obtained by restricting $\phi^{\prime\prime}$ to the grid vertex set $X:=V(D_{1})\setminus\{w_{1}\}$; the value of $\phi_{1}$ being the subgraph of $H^{\prime\prime}$ induced by $X$. We have $|X|=n^{2}\cdot n^{2}=n^{4}$.
Since there is a natural homomorphism from $H^{\prime\prime}$ to $P$, and by 20, the parameter vertices of $\phi_{1}$ span a subpath of $P$ of length at most $418n+$ 4.
As in 18, there is a subexpression $\alpha$ of $\phi_{1}$ such that $\alpha$ is the disjoint union of two smaller subexpressions $\alpha_{1}$ and $\alpha_{2}$, such that $\alpha_{1}$ contains between $\frac{1}{3}n^{4}$ and $\frac{2}{3}n^{4}$ vertices. Hence, there is a partition $(A,B)$ of $X$, such that vertices belonging to $A$ are exactly those contained within $\alpha_{1}$, and $B$ contains the remaining vertices of $X$.
Claim 21.
In the $n^{2}\times n^{2}\,$-grid $D_{1}-w_{1}$ (where $V(D_{1}-w_{1})=X$) with the partition $(A,B)$ of $X$, there is an induced matching $M$ between $A$ and $B$ of size at least $\frac{1}{3\cdot 25}n^{2}$.
Proof 5.12 (Subproof).
This is the first step of the proof of 19.
Recall that there are only $\leq 18n+4$ consecutive parameter vertices of $P$ used in $\phi_{1}$. By the pigeon-hole principle, there is a vertex $p_{i}$ of $P$ such that at least $\frac{n^{2}}{3\cdot 25}/(18n+4)=\Omega(n)$ endpoints of $M^{\prime}\subseteq M$ belonging to $A$ have $p_{i}$ as their parameter vertex. Now, in the same way as in the proof of Theorem 5.1, we conclude that the subexpression $\alpha_{1}$ requires at least $\Omega_{k}(n)$ distinct colors to define (within whole $\phi_{1}$) the required perturbation of the selected matching $M^{\prime}$. This concludes the proof.
Theorem 5.9 straightforwardly implies via Theorem 3.19 the following conclusion analogous to Corollary 5.7:
Corollary 5.13.
Let ${\cal C}$ be a graph class admitting a classical product structure. Then, ${\cal C}$ does not transduce the class of graphs ${\cal D}=\{G_{n}:n\in\mathbb{N}^{+}\}$ where $G_{n}$ is composed of $n$ disjoint copies of the pinned grid of order $n$. ∎
{toappendix}
In case connectivity of the example as in Theorem 5.9 is desirable, we add a simple modification of it. Let a pinned multigrid of order $n$ be any connected graph $G^{+}_{n}$ such that $G^{+}_{n}$ induces $n$ disjoint copies of the pinned grid of order $n$; such as, $G^{+}_{n}$ can be simply constructed from the $(n^{3}+n-1)\times n^{2}\,$-grid by suitably adding $n$ apex vertices. Theorem 5.9 then straightforwardly implies via Theorem 3.19 the following conclusion analogous to Corollary 5.7:
Corollary 5.14.
Let ${\cal D}$ be the class of graphs $G_{n}$ for $n\in\mathbb{N}^{+}$ where $G_{n}$ is composed of $n$ disjoint copies of the pinned grid of order $n$, and let ${\cal D}^{+}$ be a class containing a pinned multigrid of order $n$ for each $n\in\mathbb{N}^{+}$. If ${\cal C}$ is a graph class admitting a classical product structure, then neither of ${\cal D}$ and ${\cal D}^{+}$ is transducible from ${\cal C}$. ∎
As noted in the introduction, Gajarský, Pilipczuk and Pokrývka [DBLP:journals/corr/abs-2501-07558] give an alternative proof of Corollary 5.7 using the concept of so-called slice decomposition. It can be shown that every class of bounded ${\cal P}^{\circ}$-clique-width admits a slice decomposition, but the converse is not true in general. For example, the class ${\cal D}$ of Corollary 5.13 admits a slice decomposition.
6 Concluding Remarks
In view of Theorem 3.19 and Theorem 4.2, one comes with the following natural question. Assume ${\cal C}$ is a monadically stable class admitting a product structure, such as the class of planar graphs, and $\tau$ is a first-order transduction. Then $\tau({\cal C})$ is stable by the assumption, and up to bounded perturbations, $\tau({\cal C})$ is of bounded ${\cal P}^{\circ}$-clique-width $\leq\ell$. (Recall that ${\cal P}^{\circ}$ denotes the class of reflexive paths.) Does this mean that the perturbed graphs of $\tau({\cal C})$ always have stable $(P^{\circ},\ell)$-expressions? This may not be an easy question since, in an analogous instance, it is known that there are classes of bounded-degree planar graphs which do not admit a product structure with bounded maximum degrees of the factors [DBLP:journals/combinatorics/DujmovicJMMW24].
While we are still investigating this question, we propose:
Conjecture 6.1.
There is a function $g:\mathbb{N}\to\mathbb{N}$ such that each $({\cal Q}_{r}^{\circ},\ell)$-expressions constructed in the proof of Lemma 3.1 is $g(\ell)$-stable.
Stepping further, we define that a class ${\cal C}$ is of bounded expression-stable ${\cal H}$-clique-width if there is $k$ such that every graph of ${\cal C}$ is the value of a $k$-stable $(H,k)$-expression for $H\in{\cal H}$. A positive answer to 6.1 would, together with Theorems 3.19 and 4.2, imply that classes first-order transducible from classes admitting a product structure are exactly perturbations of those of bounded expression-stable ${\cal P}^{\circ}$-clique-width.
Previous would also imply validity of the next conjecture:
Conjecture 6.2.
The property of a graph class – to be a perturbation of a class of bounded expression-stable ${\cal P}^{\circ}$-clique-width, is preserved under taking first-order transductions.
6.2 actually formulates a strengthening of Theorem 3.19 in two ways; not only the conclusion adds expression-stable ${\cal P}^{\circ}$-clique-width, but also the assumption on the class ${\cal C}$ is weakened from admitting a product structure to having bounded expression-stable ${\cal P}^{\circ}$-clique-width. An analogous strengthening – to assume a class ${\cal C}$ of bounded expression-stable ${\cal Q}^{\circ}$-clique-width, would also be interesting in Theorem 3.16 for arbitrary bounded-degree class ${\cal Q}$.
Regarding further possibilities to strengthen Theorem 3.16, we repeat that the assumption of having a class ${\cal Q}$ of bounded maximum degree indeed is unavoidable; [DBLP:conf/mfcs/HlinenyJ24] the class of strong products of two stars is monadically independent.
We finish with two questions and a remark (Proposition 6.4) reaching slightly beyond the scope of the results here.
Firstly, we say that a class ${\cal C}$ is of bounded expression-sparse ${\cal H}$-clique-width if there is $k$ such that every graph of ${\cal C}$ is the value of an $(H,k)$-expression $\phi$ for $H\in{\cal H}$ such that the value of the deparameterization $\phi_{1}$ of $\phi$ is $K_{k,k}$-free.
Question 6.3.
Is it true that if a graph class ${\cal C}$ is of bounded ${\cal P}^{\circ}$-clique-width and ${\cal C}$ excludes bi-induced $K_{t,t}$ for some $t$, then ${\cal C}$ is also of bounded expression-sparse ${\cal P}^{\circ}$-clique-width?
Note that the value of the deparameterization $\phi_{1}$ being $K_{k,k}$-free is equivalent to this value being of bounded tree-width. Hence, in 6.3, for a class ${\cal C}$ being of bounded expression-sparse ${\cal P}^{\circ}$-clique-width is equivalent (cf. Theorem 2.4) to ${\cal C}$ admitting product structure. Being the answer to 6.3 yes, we would get kind of a ‘sparse fragment’ of Theorem 3.19 with the conclusion featuring a product structure as well.
Similarly as with the previous question about expression-stable ${\cal P}^{\circ}$-clique-width, an answer to 6.3 is not immediate (cf. [DBLP:journals/combinatorics/DujmovicJMMW24]), and we are not even sure whether this answer should be yes or no. Nevertheless, a very new result of Gajarský, Gładkowski, Jedelský, Pilipczuk and Toruńczyk [DBLP:journals/corr/abs-2505.15655] allows us to derive a ‘sparse fragment’ of Theorem 3.19 in a different way:
Proposition 6.4.
Let ${\cal C}$ be a class of graphs admitting product structure. Let $\tau$ be a first-order transduction such that $\tau({\cal C})$ is a class of simple $K_{t,t}$-free graphs for some fixed integer $t$. Then, the class $\tau({\cal C})$ “nearly admits” a product structure, meaning that there is an integer $k$ such that every graph of $\tau({\cal C})$, after deleting at most $k$ of its vertices, is a subgraph of the strong product of a path and a graph of tree-width at most $k$.
Proof 6.5.
Consider a graph $G$ and a collection ${\cal A}\subseteq 2^{V(G)}$ of vertex subsets such that, for each $A\in{\cal A}$, the induced subgraph $G[A]$ is connected. Then ${\cal A}$ is called a congestion-$c$ depth-$d$ minor model of a graph $H$ on $V(H)={\cal A}$ if: (i) every induced subgraph $G[A]$, $A\in{\cal A}$ is of radius $\leq d$, (ii) for every vertex $v\in V(G)$, at most $c$ sets of ${\cal A}$ contain $v$, and (iii) for every edge $AB\in E(H)$ (where $A,B\in{\cal A}$), the sets $A$ and $B$ touch in $G$ – meaning that $A\cap B\not=\emptyset$ or some two vertices $a\in A$ and $b\in B$ are adjacent in $G$. A graph $G_{1}$ is a congestion-$c$ depth-$d$ minor of $G$ if there exists a congestion-$c$ depth-$d$ minor model in $G$ isomorphic to $G_{1}$.
By our assumptions, the class ${\cal C}$ (as admitting product structure) is of bounded expansion and the class $\tau({\cal C})$ is $K_{t,t}$-free. Therefore, by Gajarský, Gładkowski, Jedelský, Pilipczuk and Toruńczyk [DBLP:journals/corr/abs-2505.15655], the following holds; there exists $c\in\mathbb{N}$ such that $\tau({\cal C})$ is contained in the class of congestion-$c$ depth-$c$ minors of the class ${\cal C}^{\bullet}$, where ${\cal C}^{\bullet}$ is the class obtained by adding a universal vertex to every graph of ${\cal C}$.
Let $H\in\tau({\cal C})$ which is a congestion-$c$ depth-$c$ minor of $G^{\bullet}\in{\cal C}^{\bullet}$. Let $G\subseteq G^{\bullet}$ be the graph of ${\cal C}$ from which $G^{\bullet}$ was constructed by adding a universal vertex $u$. Let $X\subseteq V(H)$ where $|X|\leq c$ be the set of vertices of $H$ whose sets in the minor model contain $u$. We are going to show that $H_{1}:=H-X$ is a subgraph of the strong product of a path and a graph of tree-width at most $k$ ($k$ independent of $H$).
First, $H_{1}$ is a congestion-$c$ depth-$c$ minor of $G\in{\cal C}$, and $G\subseteq P\boxtimes M$ where $P$ is a path and $M$ is of tree-width $\leq m$. Then, clearly, $H_{1}$ is a depth-$c$ minor of $G\boxtimes K_{c}\subseteq P\boxtimes M\boxtimes K_{c}$.
Second, Hickingbotham and Wood [DBLP:journals/siamdm/HickingbothamW24] proved (their result was actually more general) that depth-$c$ minors of graphs of the form $P\boxtimes M\boxtimes K_{c}$ are subgraphs of $P\boxtimes M^{\prime}$ where $M^{\prime}$ is of tree-width bounded only in terms of our $m$ and $c$. This concludes the proof.
Lastly, inspired by [DBLP:journals/tocl/GajarskyHOLR20], we conclude with an algorithmic question:
Question 6.6.
Let ${\cal C}$ be the class of planar graphs (or a class admitting a product structure), and $\tau$ be a first-order transduction. Can we find a first-order transduction $\tau^{\prime}$ (of ${\cal C}$) subsuming $\tau$ such that, for every graph $G\in\tau({\cal C})$ there is an efficiently computable preimage $G^{\prime}\in{\cal C}$ such that $G\in\tau^{\prime}(G^{\prime})$?